### v3

# ShaRP: A Novel Feature Importance Framework for Ranking

Abstract.
Algorithmic decisions in critical domains such as hiring, college admissions, and lending are often based on rankings. Because of the impact these decisions have on individuals, organizations, and population groups, there is a need to understand them: to help individuals improve their position in a ranking, design better ranking procedures, and check whether a procedure is legally compliant. In this paper, we present ShaRP —Shapley for Rankings and Preferences—a framework that explains the contributions of features to different aspects of a ranked outcome and is based on Shapley values. Using ShaRP , we show that even when the scoring function used by an algorithmic ranker is known and linear, the feature weights do not correspond to their Shapley value contribution. The contributions instead depend on the feature distributions and the subtle local interactions between the scoring features.
ShaRP   builds on the Quantitative Input Influence framework to compute the contributions of features for multiple —ranking specific— Quantities of Interest, including score, rank, pair-wise preference, and top-$k$. We show the results of an extensive experimental validation of ShaRP  using real and synthetic datasets. We demonstrate that feature importance can be computed efficiently, and that ShaRP  compares favorably to several prior local feature importance methods, in terms of both generality and quality of explanations. Among our results, we highlight a case study on the CS Rankings dataset. Contrary to expectation, we find that a strong track record in Systems research is much more important than AI research for placing a CS department among the top-10%.
ShaRP  is available as an open-source library at https://github.com/DataResponsibly/ShaRP and is already used in teaching.
ranking, interpretability, feature importance, Shapley values, evaluation, responsible data management, responsible AI
ccs: Information systems Data management systemsccs: Social and professional topics Socio-technical systemsccs: Human-centered computing
1. Introduction
Rankings produced by data-driven algorithmic systems now influence a myriad of socio-technical applications, as part of automated or semi-automated decision-making, and with direct consequences to people’s lives and aspirations. An algorithmic ranker, or a ranker for short, takes a database of candidates as input and produces a permutation of these candidates as output, see Figure 1 for an example. We refer to the output of a ranker as a ranked outcome or simply a ranking. As an alternative to the full permutation, the best-ranked $k$ candidates, or the top-$k$, may be returned in rank order or as a set. In the latter case, we are dealing with a selection task, which is a special case of ranking.
Algorithmic rankers are broadly used to support decision-making in critical domains, including hiring and employment, school and college admissions, credit and lending, and, of course, college ranking. Because of the impact rankers have on individuals, organizations, and population groups, there is a need to understand them: to know whether the decisions are correct and legally compliant (auditing tasks), to help individuals improve their ranked outcomes (recourse tasks), and to design better ranking procedures (design tasks). To make progress towards these tasks, we need ways to explain and interpret ranked outcomes. In this paper, we present ShaRP  —Shapley for Rankings and Preferences—a framework that explains the contributions of features to different aspects of a ranked outcome, and that can support all these critically important tasks.
name gpa sat essay $f$ $g$
Bob 4 5 5 4.6 5
Cal 4 5 5 4.6 5
Dia 5 4 4 4.4 4
Eli 4 5 3 4.2 3
Fay 5 4 3 4.2 3
Kat 5 4 2 4.0 2
Leo 4 4 3 3.8 3
Osi 3 3 3 3.0 3
$r_{\mathcal{D},f}$
Bob
Cal
Dia
Eli
Fay
Kat
Leo
Osi
$r_{\mathcal{D},g}$
Bob
Cal
Dia
Eli
Fay
Leo
Osi
Kat
Figure 1. (a) Dataset $\mathcal{D}$ of college applicants, scored on $gpa$, $sat$, and $essay$. (b) Ranking $r_{\mathcal{D},f}$ of $\mathcal{D}$ on $f=0.4\times gpa+0.4\times sat+0.2\times essay$; the highlighted top-4 candidates will be interviewed and potentially admitted. (c) Ranking $r_{\mathcal{D},g}$ on $g=1.0\times essay$; the top-$4$ coincides with that of $r_{\mathcal{D},f}$, signifying that $essay$ has the highest importance for $f$, despite carrying the lowest weight in the scoring function.
There are two types of rankers: score-based and learned. In score-based ranking, a given set of candidates is sorted on a score, which is typically computed using a simple formula, such as a weighted sum of attribute values (Zehlike et al., 2022a). In supervised learning-to-rank, a preference-enriched set of candidates is used to train a model that predicts rankings of unseen candidates (Li, 2014). In this paper, we focus on score-based rankers, which are often preferred in critical domains based on the premise that they are easier to design, understand, and justify than learning-to-rank models (Berger et al., 2019). Score-based rankers are a prominent example of the so-called “interpretable models” (Rudin, 2019): the scoring function, such as $Y_{1}=0.4\times gpa+0.4\times sat+0.2\times essay$ in a college admissions scenario, is based on a (normative) a priori understanding of what makes for a good candidate. And yet, despite being syntactically “interpretable,” a score-based ranker may not be “explainable,” in that, the designer of the ranker or the decision-maker who uses it may be unable to accurately predict or understand its output ((Miller, 2019; Molnar, 2020)). We illustrate this with an example.
Example 0.
Consider a dataset $\mathcal{D}$ of college applicants in Figure 1, with scoring features $gpa$, $sat$, and $essay$. Very different scoring functions $f=0.4\times gpa+0.4\times sat+0.2\times essay$ and $g=1.0\times essay$ induce very similar rankings $r_{\mathcal{D},f}$ and $r_{\mathcal{D},g}$, with the same top-$4$ items appearing in the same order, apparently because $essay$ is the feature that is best able to discriminate between the top-$4$ and the rest, and that determines the relative order among the top-$4$.
This example shows that the weight of a scoring feature may not be indicative of its impact on the ranked outcome. Intuitively, this happens because rankings are, by definition, relative, while feature values, and scores computed from them, are absolute. That is, knowing the score of an item says little about the position of that item in the ranking relative to other items.
There has been substantial work on quantifying feature importance for classification ((Covert et al., 2020; Datta et al., 2016; Guidotti et al., 2019; Lundberg and Lee, 2017; Mohler et al., 2020; Ribeiro et al., 2016b; Strumbelj and Kononenko, 2010)), and one may be tempted to think that these methods can be reused to explain ranking. Unfortunately, this is not the case, because in classification and regression, predictions are being made independently for each item. For example, in binary classification, if all points happen to fall on the positive side of the decision boundary, then all will receive the positive outcome. In contrast, in selection and ranking, an outcome for some item v is not independent of the outcomes for items $\mathcal{D}\setminus\{\textbf{v}\}$. That is, only one item can occupy a specific rank position, and exactly $k$ items can appear at the top-$k$.
In summary, ranking is intrinsically different from classification or regression, and it has been treated differently from classification and regression in predictive contexts, as evidenced by the rich literature on learning-to-rank (Li, 2014). Fairness in ranking has also been treated differently from fairness in classification (Zehlike et al., 2022a, b). Similarly, interpretability methods have to be customized for ranking, as we do in this work. The key insight then is that the impact of the scoring features on the ranking must be measured in a way that captures the lack of independence between per-item outcomes.
1.1. Summary of Our Approach
In this paper, we focus on local feature-based explanations of ranked outcomes. By “local,” we mean that we will explain the outcome for a particular item or a pair of items, although we may sometimes aggregate explanations across items to provide a summary. By “ranked outcome” we meant an item’s score, rank, its presence or absence among the top-$k$ (for some value of $k$), and the relative order among a pair of items. For consistency with prior literature, we will refer to different kinds of outcomes as “quantities of interest”, or QoIs for short (Datta et al., 2016). Finally, by “feature-based” we mean that we will quantify the impact of a feature’s value on a QoI.
We build on the conceptual Quantitative Input Influence (QII) framework by Datta et al. (2016) to quantify feature importance in predictive classification based on Shapley values ((Shapley et al., 1953)), a canonical method of dividing revenue in cooperative games. QII explains the influence of an item’s features (players) on the classification outcome for that item (cooperative game). We use QII as the starting point because this framework flexibly and naturally accommodates new quantities of interest. We augment QII with several Quantities of Interest (QoI) that are appropriate for ranking, including score, rank, pair-wise preference, and top-$k$.
We now preview some of the local feature-based explanations of ranked outcomes generated with ShaRP  over CS Rankings, a real dataset that ranks 189 Computer Science departments in the US based on a normalized publication count of the faculty across 4 research areas: AI, Systems, Theory, and Interdisciplinary (Berger, 2023) (see Section 7.1 for a detailed description of the dataset and the ranker). We will definite the QoIs that correspond to these ranked outcomes formally in Section 4 and will explain how ShaRP  computes them in Section 5. Here, our goal is to highlight some interesting insights about the data that these explanations can surface and, further, to show that these insights can be very different depending on what kind of ranked outcome is being explained.
(a) Score QoI: Interdisciplinar is the most important feature, followed by AI and Systems, while Theory negatively impacts the score.
(b) Rank QoI: Systems is the most important feature, followed by AI and Interdisciplinary. Theory is minimally but positively impacting the rank.
Figure 2. Feature importance for Texas A&M in CS Rankings.
Example 0.
Figure 2 presents a feature-based explanation for Texas A&M in CS Rankings, where it has a score of 3.941 and is ranked in position 34. Figures 2a and 2b show feature importance for the score and rank QoIs, respectively. Intuitively, the score QoI quantifies the impact of the features on the score, as compared to the expected (average) score. The average score in CS Ranking is 2.722, shown as $E[f(X)]=2.722$ in Figure 2a. Features shown in red in that figure have a positive impact: they increase the score of Texas A&M by the increment that is shown inside the arrow (e.g., +0.49 for “Inter.”), as compared to the average score. Features shown in blue have a negative impact; they decrease the score compared to the average (e.g.,-0.11 for “Theory”). The sum of feature impact values corresponds to the difference between the average score of 2.722 and Texas A&M’s score of 3.941, up to rounding.
Similarly, rank QoI quantifies the impact of the features on the rank, as compared to the expected rank (94.5 out of 189 in CS Ranking), but with an important notational difference: a higher score but a lower rank is desirable. For this reason, feature importance for “Systems” — the most impactful feature when it comes to improving the rank of Texas A&M — has a negative sign, but it is still shown in red, for consistency. Finally, note that the absolute value of feature importance is higher for rank than for score (e.g.,-20.78 for “Systems”) because it quantifies the actual displacement in the ranking.
According to Figure 2a, the feature “Interdisciplinary” is the most (positively) impactful when it comes to improving the score of Texas A&M, followed by “AI” and then by “Systems”, while “Theory” contributes negatively. In contrast, all features contribute to improving the rank of Texas A&M, with “Systems” being the most impactful.
Example 2 demonstrates that feature-based explanations are different for different outcomes. If the goal is to improve the score, then Texas A&M should invest in interdisciplinary and AI research (features with highest scores in Figure 2a). However, if the goal is to directly improve the rank, then it’s better to prioritize systems research. Why are the score and rank QoI explanations different? The reason is that a large change in score may not result in a comparable change in rank (and vice versa). On a related note, average score is influenced by outliers, while rank is not. In fact, over 70% of the departments in CS Rankings have scores below the mean, and, as a result, when using score QoI, many or all of their features appear to be contributing negatively. (Figure 2b shows the score vs. rank distribution for this dataset.)
Example 0.
As our final example that illustrates two other QoIs supported by ShaRP , consider Figure 3a that presents a summary view of feature contributions to the top-$k$ in CS Rankings. In this figure, the importance of each of the four features is summarized for the first 10% of the departments, followed by the next 10%, and so on. Perhaps unexpectedly, we observe that a strong publication record in systems is the most important feature for placing a CS department among the top-10, significantly ahead of AI.
Figures 3b3d analyzes the relative order among three departments that appear among the top-10: Georgia Tech (rank 5), Stanford (rank 6) and UMich (rank 7). We use visualizations for the pairwise QoI to explain that Georgia Tech ranks better than Stanford because of its relative strength in systems, and, further, that Stanford ranks better than UMich despite Stanford’s relative weakness in systems.
(a) Feature contribution to the top-$k$ QoI, for $k=10\%$. Systems is the most important feature, followed by Interdisciplinary and AI.
Institution AI Systems Theory Inter. Rank
Georgia Tech 28.5 7.8 6.9 10.2 5
Stanford 36.7 5.4 13.3 11.5 6
UMich 30.4 9.0 9.3 5.9 7
(b) Feature values and rank of three highly ranked departments: Georgia Tech, Stanford, and UMich.
(c) Pairwise QoI: Georgia Tech ranks higher than Stanford because of its relative strength in Systems.
(d) Pairwise QoI: Stanford ranks higher than UMich despite Stanford’s relative weakness in Systems.
Figure 3. Feature importance for the top-$k$ QoI (i.e., selection) for CS Rankings in 3a, with further analysis of the relative orders among two pairs of departments in 3c and 3d.
1.2. Contributions
Our first contribution is a formalization of several natural quantities of interest (QoI) for ranking, expressing feature contributions to an item’s score or rank, to its presence or absence at the top-$k$, and to the relative order between a pair of items. To define the QoIs, we build on the conceptual Quantitative Input Influence (QII) framework by Datta et al. (2016) and extend it to quantify the influence of features on the ranked outcome. QII quantifies feature importance based on Shapley values ((Shapley et al., 1953)), a canonical method of dividing revenue in cooperative games. Like other Shapley value methods, QII explains the influence of an item’s features (players) on the classification outcome for that item (cooperative game). Unlike other Shapley value methods, QII does not necessarily use the score of the model being explained as the payoff but allows for the definition of a variety of payoff functions, termed Quantities of Interest (QoIs). This flexibility is essential for explaining ranking, because of the need to capture the lack of independence between per-item outcomes.
QII is a conceptual framework and only a preliminary implementation is publicly available. Our second contribution is a robust and extensible open-source ShaRP  library that implements QII, making this powerful framework — including also its functionality for explaining classification outcomes — available to the community. ShaRP  is available as an open-source library at https://github.com/DataResponsibly/ShaRP and we already use it to teach interpretability to computer science and data science students. Our implementation includes approximate QoI computation to achieve interactive performance while retaining fidelity and consistency.
Our third contribution is an extensive experimental evaluation of ShaRP . We define metrics for evaluating ranking explanability methods, and present results of qualitative analysis of ShaRP  with numerous real and synthetic datasets, showing that feature importance for all QoIs depends on the properties of the data distribution even when the scoring function remains fixed. Further, we show that the QoIs we define bring interesting — and complementary — insights about the data that cannot be captured simply by the relationship between the scores and the ranks. We also present results of a large-scale quantitative evaluation of ShaRP , demonstrating the scalability of approximations, and comparing it to other local feature importance methods.
2. Related Work
Local feature-based explanations
Ribeiro et al. (2016b) proposed LIME, a feature-based method that explains a complex classification model locally using a class of more interpretable models such as sparse linear models.  Lundberg and Lee (2017) proposed SHAP, a feature-based method that uses Shapley values to explain a large variety of classification and regression models, and that generalizes LIME. Both LIME and SHAP have been implemented in broadly used software libraries, and they can be used to explain the score of an item (score QoI in our terminology). We compare with both SHAP and LIME experimentally in Section 7.4.
Feature-based explanations for ranking
Yang et al. (2018) proposed a “nutritional label” for score-based rankers that includes two visual widgets that explain the global ranking methodology and outcome: “Recipe” shows scoring feature weights and “Ingredients” shows the features (scoring or not) that have the strongest Spearman’s rank correlation with the score. The authors observed that the weight of a feature in the “Recipe” rarely corresponds to how well it correlates with the score. This work generates global feature-based explanations, showing which features matter more or less for the ranked outcome for all items. Another very recent line of work, Heuss et al. (2024), adapts SHAP (Lundberg and Lee, 2017) for learning-to-rank, to compute global feature-based explanations. In contrast, we focus on local explanations that show feature importance for the outcome of a specific item or a pair of items.
Gale and Marian (2020) developed “participation metrics” for score-based rankers that measure the contributions of scoring features to whether an item is among the top-$k$. Their most important metric, “weighted participation,” attributes an item’s presence among the top-$k$ to the values of the scoring features and their scoring function weights, to the distance of a feature value from the maximum possible values, and to whether the value exceeds the $k^{th}$ highest value for the feature. This quantity is then aggregated across all items at the top-$k$, and outputs one participation value per feature. The main difference between our method and Gale and Marian (2020) is that we generate local — per item — explanations, while their explanations are group-wise, for a specific group — the top-$k$. Our top-$k$ QoI is similar to their weighted participation, however, our focus is on local — per item — explanations, which we can then (optionally) aggregate across the top-$k$. As a result, we cannot compare with Gale and Marian (2020) directly.
Yuan and Dasgupta (2023) developed an interface for sensitivity analysis of ranked synthetic datasets with two normally distributed features, and with a scoring function with linear feature weights. They quantify the importance of each feature by subtracting its value from the mean value of that feature. Given their distributional assumptions and assumptions about the scoring function, this corresponds to an approximation of Shapley values of the features (Lundberg and Lee, 2017). They also propose two alternative methods to quantify feature contributions, calculated using linear weights and the score/rank of the item. We re-implemented and substantially extend their method to work with more than 2 features with arbitrary feature distributions, and with arbitrary scoring functions. This allowed us to compare with their method experimentally in Section 7.4.
Anahideh and Mohabbati-Kalejahi (2022) provide local feature-based explanations using multiple methods, including Shapley values computed with SHAP (Lundberg and Lee, 2017), and apply these methods to the neighboring items ranked around the item being explained. They justify this approach by observing that feature importance changes when different size ranking neighborhoods are considered. Like Anahideh and Mohabbati-Kalejahi (2022), we find that feature importance differs significantly per rank stratum. However, countering the assumptions of  Anahideh and Mohabbati-Kalejahi (2022), we show that slight changes in feature values can move an item significantly in the ranks, far beyond the local neighborhood. We will conduct an experimental comparison with their method in Section 7.4.
Finally,  Hu et al. (2022) developed PrefSHAP, a method for explaining pairwise preference data from learned rankers. They calculate the Shapley value contributions for pairwise preferences on a new quantity they propose: They create artificial items from the pair of items simultaneously for each coalition, and then evaluate how the preference changes. While this quantity can be applied to any preferential function, they introduce PrefSHAP specifically for functions that use the generalized preferential kernel defined in their own prior work (Chau et al., 2022) for running time efficiency. They compare their results with an adaptation of SHAP for preference models and show that this adaptation does not capture the contributions properly. We also argue that ranking-specific quantities of interest are needed to explain rankings and introduce a broader set of QoIs (score, rank, top-$k$, and pairwise). We do not conduct an experimental comparison between our pairwise QoI and PrefSHAP Hu et al. (2022) because our focus is on preferences that are induced by score-based rankers, while they work only with a specific kind of learned models.
In summary, we share motivation with these lines of work but take a leap, describing the first comprehensive Shapley-value-based framework for selection and ranking.
3. Preliminaries and Notation
Ranking.
Let $\mathcal{A}$ denote an ordered collection of features (equiv. attributes), and let $\mathcal{D}_{\mathit{}}$ denote a set of items (equiv. points or candidates). An item $\textbf{v}=(v_{1},\dots,v_{d})\in\mathbb{R}^{d}$ assigns values to $|\mathcal{A}|=d$ features, and may additionally be associated with a score. Score-based rankers use a scoring function $f(\textbf{v})$ to compute the score of v. For example, using $f_{1}(\textbf{v})=0.4\times gpa+0.4\times sat+0.2\times essay$, we compute $f(\texttt{Bob})=4.6$ and $f(\texttt{Leo})=3.8$.
A ranking $r_{\mathcal{D}}$ is a permutation over the items in $\mathcal{D}_{\mathit{}}$. Letting $n=\left|\mathcal{D}_{\mathit{}}\right|$, we denote by $r_{\mathcal{D}}=\langle\textbf{v}_{1},\ldots,\textbf{v}_{n}\rangle$ a ranking that places item $\textbf{v}_{i}$ at rank $i$. We denote by $r_{\mathcal{D}}(i)$ the item at rank $i$, and by $r_{\mathcal{D}}^{-1}(\textbf{v})$ the rank of item v in $r_{\mathcal{D}}$. In score-based ranking, we are interested in rankings induced by some scoring function $f$. We denote these rankings $r_{\mathcal{D},f}$. For example, in Figure 1(b), $r_{\mathcal{D},f}(1)=\texttt{Bob}$, $r_{\mathcal{D},f}^{-1}(\texttt{Leo})=7$. We assume that $r_{\mathcal{D},f}^{-1}(\textbf{v}_{1})<r_{\mathcal{D},f}^{-1}(\textbf{v}_{2})<\dots<r_{\mathcal{D},f}^{-1}(\textbf{v}_{n})$, where smaller rank means better position in the ranking.
[…]
Our goal is to explain the importance of features $\mathcal{A}$ to the ranking $r_{\mathcal{D},f}$. We will do so using Shapley values (Shapley et al., 1953).
Shapley values
For a set $\mathcal{N}$ of $n$ players, and a value function $f$ that assigns a profit to any subset (or coalition) $\mathcal{S}$ of players, $f:2^{n}\rightarrow\mathbb{R}$, where $f(\emptyset)=0$, the Shapley value of player $i$ is:
(1) $\phi_{i}(f)=\sum_{\mathcal{S}}\frac{|\mathcal{S}|!(n-|\mathcal{S}|-1)!}{n!}(f(\mathcal{S}\cup\{i\})-f(\mathcal{S}))$
We will use Shapley values to explain ranked outcomes using the set of features $\mathcal{A}$ as the players, and the outcome (or the quantity of interest, QoI) as the payoff function. In addition to the definition of players and the payoff function, Shapley values require the quantification of the payoff over a subset of the players. This, in turn, requires some notion of estimating the payoff over a subset of the features. Consequently, for any Shapley value implementation, a method of feature removal or masking is required (Covert et al., 2021).
A common method (e.g., used in SHAP (Lundberg and Lee, 2017; Covert et al., 2021)), for a coalition (subset of features) $\mathcal{S}\subseteq\mathcal{A}$, is to marginalize out the features not in the coalition $\mathcal{A}\setminus\mathcal{S}$ and draw values from the marginal distributions of the subset of features in $\mathcal{S}$ jointly. Another alternative (e.g., used in QII (Datta et al., 2016)) is to draw values of each feature in $\mathcal{S}$ independently from its marginal distribution.
More specifically, let $\textbf{v}_{\mathcal{S}}$ denote a projection of v onto $\mathcal{S}$. In the example in Figure 1, $(\texttt{Bob},\texttt{4},\texttt{5},\texttt{2})_{\{name,gpa\}}=(\texttt{Bob},\texttt{4})$. We define a random variable U that draws values from the marginal distributions of the subset of features in $\mathcal{S}$. Let U $=\langle\textbf{u}_{1},\dots,\textbf{u}_{m}\rangle$ denote a vector of $m$ items sampled from $\mathcal{D}$ using this method. For a subset of features $\mathcal{S}\in\mathcal{A}$, let $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}=\langle\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{1})_{\mathcal{S}},\dots,\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{m})_{\mathcal{S}}\rangle$ denote a vector of items, in which each $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{i})_{\mathcal{S}}$ takes on the values of the features in $\mathcal{S}$ from $\textbf{u}_{i}$, and the values of the remaining features $\mathcal{A}\setminus\mathcal{S}$ from v. We calculate Shapley values using this set of features $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$, note that if $m=|\mathcal{D}-1|$ we use the entire dataset $\mathcal{D}\setminus\textbf{v}$ to calculate the exact Shapley values.
Shapley values satisfy several natural axioms, including efficiency, symmetry, dummy, and additivity (Shapley et al., 1953), with additional useful properties, such as monotonicity, following from these axioms(Lundberg and Lee, 2017). Efficiency states that the sum of the contributions of all features for item v equals the QoI: $\sum_{i\in\mathcal{A}}\varphi_{i}(f,\textbf{v})=f(\textbf{v})$. Using this property, a Shapley-value-based explanation can be used to reconstruct the outcome, potentially leading to the explanation being more actionable, as we argued in Example 2. We will use the efficiency property to define the fidelity metric for comparing explanations (Section 6).
4. Quantities of Interest for Ranking
The first contribution of our work is that we define QoIs that are appropriate for ranked outcomes. In addition to the expected score, we also introduce rank, top-$k$, and pairwise QoI.
Score QoI
The Shapley value function for the score QoI is:
(2) $QoI_{f,\textbf{v}}(S)=\mathop{\mathbb{E}}_{\textbf{U}_{\mathcal{S}}}[f(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
This QoI captures the impact of an item’s features on its score. This is the QoI used by the popular feature-based explanation methods such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016a). To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the score over a random variable $\textbf{U}_{\mathcal{S}}$ that draws values from the marginal distributions of the set of features in $\mathcal{S}$.
Rank QoI
The Shapley value payoff function for the rank QoI is:
(3) $QoI_{f,\textbf{v},\mathcal{D}}(S)=\mathop{\mathbb{E}}_{\textbf{U}_{\mathcal{S}}}[r^{-1}_{\mathcal{D}^{\prime},f}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. This QoI evaluates the impact of an item’s features on its rank. To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the rank over a random variable $\textbf{U}_{\mathcal{S}}$ that draws values from the marginal distributions of the set of features in $\mathcal{S}$.
Top-$k$ QoI
The Shapley value payoff function to quantify the impact of an item’s features on its presence or absence among the top-$k$ is stated similarly as rank QoI:
(4) $QoI_{f,\textbf{v},\mathcal{D}}(S)=\mathbb{E}_{\textbf{U}_{\mathcal{S}}}[\mathbbm{1}_{r_{\mathcal{D}^{\prime},f}(1\ldots k)}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. The difference with rank QoI (Equation 3) is that here we compute the expectation over the indicator function that returns 1 if an items’ rank is at most $k$ and 0 otherwise.
Pairwise QoI
The Shapley value payoff function for the pairwise explanation of the difference in the ranks between items v and u is defined on $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}$ instead of $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$:
(5) $QoI_{f,\textbf{v},\mathcal{D}}(S)=r^{-1}_{\mathcal{D}^{\prime},f}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}})$
Using this function, we quantify the contribution of each feature of v to the relative order (preference) between the two items.
5. Implementation of ShaRP
We implement the QoIs of Section 4 in the ShaRP  library. [^1] ShaRP  is available as an open-source library at https://github.com/DataResponsibly/ShaRP.  ShaRP  is implemented in Python, follows an API structure similar to scikit-learn (Buitinck et al., 2013), and is parallelized. The library can be used both to compute exact feature importance values and to approximate them to improve running times. It also implements several explanation presentation methods that are appropriate for different QoIs.
5.1. Implementation of QoIs for ranking
Algorithm 1 presents the computation of feature importance — as a vector of Shapley values $\phi(\textbf{v})$ — to QoIs that can be specified for a single item v, independently of other items. The algorithm relies on black-box access to the model that generates the outcomes (i.e., specifying an input to the model and observing the output).
Algorithm 1 takes as input a dataset $\mathcal{D}$, an item v for which the explanation is generated, the number of samples to draw $m$, and the QoI function $\iota()$ based on which to quantify feature importance.
Note that parameter $m$ is user-defined and controls the quality of approximation. Passing in the full set of items $\mathcal{D}$ and setting $m=|\mathcal{D}-1|$ results in exact computation of the Shapley values (i.e., the item v’s features are quantified against every other item in $\mathcal{D}$). Similarly, in line 3, we can restrict the maximum size of $\mathcal{S}$ and reduce the number of coalitions we use. We will provide an empirical analysis of the impact of $m$ and of maximum coalition size on performance in Section 7.2.2.
For every feature $i\in\mathcal{A}$, the algorithm computes every coalition of the remaining features $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$. For each coalition $\mathcal{S}$, it draws $m$ samples from $\mathcal{D}$. Two vectors of items are created based on this sample: $\textbf{U}_{1}$, in which feature values of the coalition vary as in U and the remaining features are fixed to their values in v; and $\textbf{U}_{2}$, in which features in the coalition and the feature of interest $i$ vary as in U, and the remaining features are fixed to their values of v. The importance of the coalition, $\phi_{i_{\mathcal{S}}}(\textbf{v})$, is computed by invoking the function $\iota()$ that computes the difference in QoI between the two vectors of items $\textbf{U}_{1}$ and $\textbf{U}_{2}$. Following (Datta et al., 2016), we define $\iota_{QoI}(\textbf{U}_{1},\textbf{U}_{2})=QoI(\textbf{U}_{1})-QoI(\textbf{U}_{2})$. This value is added to the contribution of feature $i$, $\phi_{i}(\textbf{v})$ weighted by the number of coalitions of this size (${d-1\choose|S|}$, and the total number of possible coalition sizes $d$.
Algorithm 1 Feature importance for per-item outcomes
0:  Dataset $\mathcal{D}$, item v, number of samples $m$, $\iota()$
1:  $\phi(\textbf{v})=\langle 0,\ldots,0\rangle$
2:  for $i\in\mathcal{A}$ do
3:     for  $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$ do
[…]
5:        $\textbf{U}_{1}=\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$
6:        $\textbf{U}_{2}=\textbf{v}_{\mathcal{A}\setminus\{\mathcal{S}\cup i\}}\textbf{U}_{\mathcal{S}\cup i}$
7:        $\phi_{i_{\mathcal{S}}}(\textbf{v})=\iota(\textbf{U}_{1},\textbf{U}_{2})$
8:        $\phi_{i}(\textbf{v})=\phi_{i}(\textbf{v})+\frac{1}{d}\frac{1}{{d-1\choose|S|}}\phi_{i_{\mathcal{S}}}(\textbf{v})$
9:     end for
10:  end for
11:  return  $\phi(\textbf{v})$, the Shapley values v’s features
We use Algorithm 1 as a starting point for computing feature importance for ranked outcomes.
Score QoI
Recall that $f$ denotes the scoring function. We overload notation and denote by $f(\textbf{U})$ the computation of the score for each item in U, and return these scores as a vector. Since $\textbf{U}_{1}$ and $\textbf{U}_{2}$ are vectors of items constructed from the same vector U, they are of the same size. To quantify feature importance to an item’s score we use the $\iota()$ function and the QoI defined in the previous section, to take the mean of the (per-element) difference of $f(\textbf{U}_{1})$ and $f(\textbf{U}_{2})$.
Rank QoI
An item’s rank is computed with respect to all other items in the sample. For this reason, to quantify the impact of v’s features on its rank, we must compute the ranking of each intervened-upon instantiation of v in $\mathcal{D^{\prime}}$. This adds two steps to calculating this QoI compared to the score QoI. The item we are explaining needs to be removed from $\mathcal{D^{\prime}}$, and the score of each item $\textbf{u}_{i}\in\textbf{U}_{1}$ (and equivalently $\textbf{u}_{j}\in\textbf{U}_{2}$) needs to be compared to the scores of all items in $\mathcal{D^{\prime}}$.
Algorithm 2 $\iota_{Rank}$
0:  Dataset $\mathcal{D}$, scoring function $f$, item v, $\textbf{U}_{1}$, $\textbf{U}_{2}$, number of samples $m$
1:  $\phi=0$
2:  for $i\in\{1,\dots,m\}$ do
3:     $\textbf{u}_{1}=\textbf{U}_{1}(i)$
4:     $\textbf{u}_{2}=\textbf{U}_{2}(i)$
5:     $\mathcal{D}_{1}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{1}\}$
6:     $\mathcal{D}_{2}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{2}\}$
7:     $\phi=\phi+r^{-1}_{\mathcal{D}_{2},f}(\textbf{u}_{2})-r^{-1}_{\mathcal{D}_{1},f}(\textbf{u}_{1})$
8:  end for
9:  return  $\phi/|\textbf{U}_{1}|$
The computation of $\iota_{Rank}$ is summarized in Algorithm 2. We consider each a pair of corresponding (i.e., at the same position in their respective vector) samples $\textbf{u}_{1}\in\textbf{U}_{1}$ and $\textbf{u}_{2}\in\textbf{U}_{2}$, and compute $\mathcal{D}_{1}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{1}\}$ and $\mathcal{D}_{2}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{2}\}$. Next, we rank $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ on $f$, and compute the difference in ranks $\iota_{rank}=r^{-1}_{\mathcal{D}_{2},f}(\textbf{u}_{2})-r^{-1}_{\mathcal{D}_{1},f}(\textbf{u}_{1})$. Note that we are subtracting the rank of $\textbf{u}_{1}$ from the rank of $\textbf{u}_{2}$ since a lower rank corresponds to a higher (better) position in the ranking. We return the average (expected) rank difference as $\iota$.
Top-$k$ QoI
To compute feature importance that explains whether an item appears at the top-$k$, for some given $k$, we use a similar method as for rank QoI. The difference is that rather than computing the difference in rank positions for a given pair of items $\textbf{u}_{1}$ and $\textbf{u}_{2}$, we instead check whether one, both, or neither of them are at the top-$k$. As in Algorithm 2, we work with $\mathcal{D}_{1}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{1}\}$ and $\mathcal{D}_{2}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{2}\}$ for each sample. We increase the contribution to $\phi$ by 1 if only $\textbf{u}_{1}$ is in the top-$k$, and decrease it by 1 if only $\textbf{u}_{2}$ is in the top-$k$. We omit pseudocode due to space constraints.
Pairwise QoI
[…]
0:  items v and u, $\iota()$
1:  $\phi(\textbf{u}\succ\textbf{v})=\langle 0,\ldots,0\rangle$
2:  for $i\in\mathcal{A}$ do
3:     for $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$ do
4:        $\textbf{U}_{1}=\langle\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}\rangle$
5:        $\textbf{U}_{2}=\langle\textbf{v}_{\mathcal{A}\setminus\{\mathcal{S}\cup i\}}\textbf{u}_{\mathcal{S}\cup i}\rangle$
6:        $\phi_{i_{\mathcal{S}}}(\textbf{u}\succ\textbf{v})=\iota(\textbf{U}_{1},\textbf{U}_{2})$
7:        $\phi_{i}(\textbf{u}\succ\textbf{v})=\phi_{i}(\textbf{u}\succ\textbf{v})+\frac{1}{d}\frac{1}{{d-1\choose|S|}}\phi_{i_{\mathcal{S}}}(\textbf{u}\succ\textbf{v})$
8:     end for
9:  end for
10:  return  $\phi(\textbf{u}\succ\textbf{v})$ that explain why u is ranked higher than v
5.2. Visualizing Feature Importance
We use three methods for visualizing experimental results.
The first, used in Figure 2, is a waterfall visualization that presents feature importance for a single item (as in (Lundberg and Lee, 2017)).
The second visualization, used in Figures 3a,  9,  10,  12, and 11, is a box-and-whiskers plot that aggregates local feature importance across ranking strata, each covering 10% of the range. The box and whiskers plots allow us to see both the median contribution of each feature and the variance in that feature’s contribution across the items in the stratum. We choose to visualize the contributions in different strata, instead of one global feature importance plot because the contributions differ per ranking stratum. Due to the ranking process, the items in each stratum have different ranges of score values. This implies that each stratum contains items with different ranges of feature values. As a result, the feature contributions per stratum can be significantly different, see Figure 9 for an example. We can further characterize the ranges of the features of each stratum if we take into account the feature distributions and the scoring function, see discussion in Section 7.3.2 for an example.
The third visualization, used in Figures 3c and 3d, is a bar graph that presents feature contributions for pairwise QoI, created from the point of view of the first item of the pair. Note that, due to the efficiency property, defined in Section 3, feature contributions add up to the difference in ranks between the two items.
6. Metrics
Recall that our focus is on explanation methods that return a numerical vector of feature attributions explaining the outcome for a given item. We can assess pair-wise explanation agreement by comparing the feature vectors of a pair of explanations. Furthermore, we can use an explanation to compute the outcome for the item being explained (e.g., its rank), and compare it to the actual observed outcome for that item. This allows us to assess fidelity of an explanation. Below, we describe explanation agreement and fidelity metrics and also explain how these primitives can be aggregated to assess sensitivity and fidelity of an explanation method, and to quantify inter-method explanation agreement.
Notation
In Section 3, we have been using $\phi(\textbf{v})$ to represent the vector of feature weights, computed using Shapley values. We generalize our definition here to $g(\textbf{v})$ to represent the output of any feature-based explanation method $g$, regardless of whether it consists of Shapley values or of some other quantification of feature importance, with weights. For all methods we consider, $g(\textbf{v})$ is a vector of numerical contributions of each feature towards the outcome for item v.
6.1. Fidelity Metrics
Explanation Fidelity
A useful property of feature-based explanations is that the actual outcome can be computed from them. For Shapley-value-based explanations, this follows from the efficiency property of Shapley values, see Section 3. Fidelity measures how well the explanation $g(\textbf{v})$ matches the model prediction $f(\textbf{v})$ being explained, see (Chowdhury et al., 2023; Molnar, 2020). SHAP and LIME explanations can be used to compute an item’s score (score QoI in our terminology), with feature importance indicating the displacement due to that feature from the mean score, either positively or negatively. ShaRP  explanations can be used to compute the outcome for all supported QoIs, including score, rank, top-$k$, and pairwise.
For QoIs that concern a single item, namely, score, rank, and top-$k$, we compute fidelity of explanation $g$ for item v as:
(6) $F(g,\textbf{v},\mathsf{QoI}())=1-\frac{1}{Z}\left|{\mathsf{QoI}(\textbf{v})-\sum_{i=1}^{d}g(i,\textbf{v})}\right|$
Here, $\mathsf{QoI}(\textbf{v})$ returns the value of the quantity of interest (i.e., the outcome being explained by $g$), such as v’s score, rank, or presence in the top-$k$, while $g(i,\textbf{v})$ is the contribution of the i-th feature of $v$. Finally, $Z$ is the normalizer set to the maximum distance between a pair of outcomes for the given dataset $\mathcal{D}$ and ranker $f$ (omitted to simplify notation), and for the specified QoI. Note that, for pairwise QoI, fidelity $F(g,\textbf{u}\succ\textbf{v})=1$ if $u$ is ranked higher than $u$ and if $g$ predicts that relative order among the items, and is 00 otherwise.
Consider, for example, the explanation of Texas A&M University’s $rank=34$ in CS Rankings, presented as a waterfall plot in Figure 2b. The sum of feature weights $-20.78-19.52-18.71-1.95=-60.96$ captures the displacement of Texas A&M University in the ranking relative to the middle of the ranked list (position 94.5 out of 189), up to rounding: $94.5-60.96=33.54$. This explanation has near-perfect fidelity $0.9981-\frac{0.16}{189}=$ 0.998. We use the length of the ranked list $Z=189$ as the normalizer for rank QoI.
Method Fidelity
We aggregate per-item fidelity (per Equation 6) to quantify the fidelity of an explanation method as:
(7) $F(g,\mathcal{D})=\mathbb{E}_{\textbf{v}\in\mathcal{D}}F(g,\textbf{v})$
For pairwise QoI, we compute $F(g,\mathcal{D})$ as the expectation of $F(g,\textbf{u}\succ\textbf{v})$ over all pairs of distinct items $\textbf{u},\textbf{v}\in\mathcal{D}$.
6.2. Agreement Metrics
Explanation Agreement
When comparing explanation methods, we may be interested in knowing how similar their explanations are for the same item. Alternatively, when analyzing an explanation method, we may want to know how similar its explanations are for some pair of items (e.g., those that are similar in feature space, or that have similar outcomes, or both).
We define three flavors of explanation agreement, based on (1) Kendall’s tau distance, (2) Jaccard distance of the top-$2$ features, and (3) Euclidean distance between the explanation vectors. For each of these distance metrics, we normalize them to the $[0,1]$ range and then transform their output so that 1 means full agreement (similarity) and 0 means full disagreement. For dataset $\mathcal{D}$ and ranker $f$, we define explanation agreement as:
(8) $A(g,q,\textbf{u},\textbf{v},\mathsf{sim}())=\mathsf{sim}(g(\textbf{u}),q(\textbf{v}))$
[…]
Method Agreement
To compute agreement for a pair of explanation methods $g$ and $q$, for a dataset $\mathcal{D}$, we compute explanations for each item using each method, compute pair-wise explanation agreement per Eq. 8, and aggregate it across $\mathcal{D}$.
(9) $A(g,q,\mathcal{D},\mathsf{sim}())=\mathbb{E}_{\textbf{v}\in\mathcal{D}}A(g,q,\textbf{v},\textbf{v},\mathsf{sim}())$
Method Sensitivity
The Sensitivity of an explanation method quantifies the similarity between explanations of similar items (Bhatt et al., 2021). We will use $\mathsf{nbr}(\mathbf{v})$ (as in “neighbor”) to refer to a function that retrieves items that are in some sense similar to $\mathbf{v}$, noting that this similarity may be based on items’ features, their outcomes for some QoI, or both. For each v, we retrieve its neighbors $\mathsf{nbr}(\textbf{v})$, compute pair-wise explanation agreement between v and each of its neighbors per Eq. 8, and aggregate this value over $\mathcal{D}$:
(10) $S(g,\mathcal{D},\mathsf{sim},\mathsf{nbr}())=\mathbb{E}_{\textbf{v}\in\mathcal{D},\textbf{u}\in\mathsf{nbr(\textbf{v}})}A(g,g,\textbf{v},\textbf{u},\mathsf{sim}())$
7. Experimental Evaluation of ShaRP
We conducted numerous experiments to showcase the usefulness and performance of ShaRP  over real and synthetic datasets, with associated score-based ranking tasks. We present efficiency analysis in Section 7.2, conduct a qualitative evaluation of ShaRP  in Section 7.3, and compare it with several other explainability methods in Section 7.4 using the metrics from Section 6. All experiments were conducted on a Linux machine with an Intel Xeon Platinum 8268 CPU 2.90GHz Processor, using 14 cores and 128GB RAM.
7.1. Datasets
We evaluate the performance of ShaRP  and compare it to other local feature importance methods, using several real and synthetic datasets, with the corresponding ranking tasks. Dataset properties, along with ranker type (score-based or learned) are summarized in Table 1 and described below. We show the relationship between score and rank for score-based ranking tasks in Figure 4.
name source # tuples # features
Tennis (ATP) (Joao Fonseca, 2024b) 86 6
CS Rankings (CSR) (Berger, 2023) 189 5
Times Higher Education (THE) (Joao Fonseca, 2024a) 1,397 5
Synthetic (SYN) this paper 2,000 2 or 3
Table 1. Datasets used in our experimental study.
CS Rankings (CSR) ranks 189 Computer Science departments in the US based on a normalized publication count of the faculty across 4 research areas: AI, Systems (Sys), Theory (Th), and Interdisciplinary (Int) (Venetia Pliatsika, 2023). We use publication data for 2013-2023, with the scoring function provided by CSRankings, a geometric mean of the adjusted counts per area, with # of sub-areas as exponent:
$f=\sqrt[27]{(AC_{AI}^{5}+1)(AC_{Sys}^{12}+1)(AC_{Th}^{3}+1)(AC_{Int}^{7}+1)}$
ATP Tennis (ATP) is based on publicly available 2020-2023 performance data of tennis players from the Association of Tennis Professionals (ATP) (Joao Fonseca, 2024b). We use 2022 data that includes 5 performance-related attributes of 86 players. We select 2022 because this is the year in which data for all 5 attributes is available for the highest number of players. We use the following scoring function that we recovered from the ATP site using the scores:
[…]
Times Higher Education (THE) is a dataset of worldwide university rankings (Joao Fonseca, 2024a). It contains the university name, country, and the scores assigned to the university by Times Higher Education for teaching (TEA), research (RES), citations (CIT), income (INC), and international students (INT). We use 2020 data, for consistency with Anahideh and Mohabbati-Kalejahi (Anahideh and Mohabbati-Kalejahi, 2022) who also used it in their paper, with the scoring function provided by THE:
$f=0.3\times TEA+0.3\times RES+0.3\times CIT+0.025\times INC+0.075\times INT$
(a) Tennis (ATP)
(b) CS Rankings (CSR)
(c) Times Higher Education (THE)
(d) a representative synthetic dataset
Figure 4. The relationship between an item’s score ($y$-axis) and its rank ($x$-axis) for four score-based tasks.
Synthetic datasets
We also use numerous synthetic datasets to showcase specific quantitative and qualitative aspects of local feature-based explanations and metrics, and to study specific aspects of performance. These datasets contain 2,000 tuples. In five of them, items have 2 features, $x_{1}$ and $x_{2}$, distributed according to the uniform, Gaussian, or Bernoulli distributions, with varying parameters. We experiment with both independent and correlated features. Each synthetic dataset consists of $2,000$ items. We use three linear scoring functions: $f_{1}=0.8\times x_{1}+0.2\times x_{2}$, $f_{2}=0.5\times x_{1}+0.5\times x_{2}$, and $f_{3}=0.2\times x_{1}+0.8\times x_{2}$.
To explore correlations further, we create three datasets that have three Normal features $x_{1}$, $x_{2}$, and $x_{3}$, and 2,000 items. In the first dataset, all features are independent. In the second, we draw $x_{1}$ and $x_{2}$ from the 2D Gaussian and they are negatively correlated with correlation -0.8. The third feature $x_{3}$ is independent. For the third dataset, we draw the features from the 3D Gaussian. $x_{1}$ and $x_{2}$ are negatively correlated with correlation -0.8, $x_{1}$ and $x_{3}$ are positively correlated with correlation 0.6, and $x_{2}$ and $x_{3}$ are negatively correlated with correlation -0.2. For all three datasets, we use the same scoring function $f_{4}=0.33\times x_{1}+0.33\times x_{2}+0.34\times x_{3}$.
avg. time (sec)
dataset # tuples # features score rank pair
ATP 86 6 0.004 0.026 0.004
CSR 189 5 0.002 0.022 0.003
THE 1,397 5 0.011 0.423 0.007
SYN 2,000 3 0.002 0.126 0.003
Table 2. Running time of exact computation.
7.2. Efficiency and Approximation
(a) Speedup vs. sample size, THE
(b) Speedup vs. sample size, CSR
(c) Speedup vs. sample size, ATP
(d) Speedup vs. max coalition size, THE
(e) Speedup vs. maximum coalition size, CSR
(f) Speedup vs. maximum coalition size, ATP
Figure 5. Computational time performance of approximation for Times Higher Education (THE), CSRankings (CSR), and ATP Tennis (ATP). Speedup is computed in comparison to exact computation times, reported in Table 2.
7.2.1. Running time of exact computation
By definition, when computing feature importance, Shapley values use all possible coalitions of features and all items in the dataset. We refer to this as the exact computation of local feature-based explanations. In our first experiment, we measure the running time of the exact computation for rank, score, and pairwise QoI computation for 3 of the real and 1 of the synthetic datasets in Table 1. We include only 1 synthetic dataset because all synthetic datasets have the same number of points ($m=2,000$) and at most 3 features, and differ in the structure of the correlations within the data. However, these correlations do not impact the running time of exact computation. Finally, we do not include the running time for the top-$k$ QoI because it is based closely on rank QoI in our implementation, and the running times of generating explanations for these two QoIs are indistinguishable.
Table 2 presents the results, reporting the time to generate an explanation per point, averaged over 100 points for CSR, THE, and SYN, and over 83 points (dataset size) for ATP. We observe that rank QoI is the most challenging to compute, taking 0.423 seconds per item for THE, followed by score (0.011 seconds for THE) and then by pairwise (0.007 seconds for THE).
Further, we observe that the running time of rank and score QoIs increases both with the number of items ($m$ in Algorithm 1) and with the number of features $d$. This is as expected, since exact feature importance according to score and rank QoIs depends linearly on dataset size $\mathcal{D}$ (since $m=|\mathcal{D}|$ samples are taken for exact computation) and exponentially on the number of features $d$ (there are $12^{d}-$ 1 feature coalitions).
The computation of pairwise QoI only uses two points in the dataset (the points making up the pair), and so we would expect its running time to be insensitive to $m$ but exponential in $d$. Note, however, that both rank and pairwise QoI require computing the new rank of an item, or pair of items, upon intervention as part of Shapley value computation (line 7 in Algorithm 2 for rank QoI, and line 6 in Algorithm 3 for pairwise QoI). This rank computation can be implemented in a number of ways. In our implementation, this operation is linear in dataset size $m$, which explains, for example, that the running time of pairwise QoI for THE ($m=1,397$ and $d=5$) is higher than the running time for ATP ($m=86$ and $d=6$).
(a) Fidelity vs. maximum sample size, THE
(b) Fidelity vs. maximum sample size, CSR
(c) Fidelity vs. maximum sample size, ATP
(d) Fidelity vs. max coalition size, THE
(e) Fidelity vs. maximum coalition size, CSR
(f) Fidelity vs. maximum coalition size, ATP
Figure 6. Fidelity of approximation for Times Higher Education (THE), CSRankings (CSR), and ATP Tennis (ATP), using different sample sizes and maximum coalition sizes.
(a) Agreement of ShaRP across different sample sizes using maximum coalition size for rank QoI
(b) Agreement of ShaRP across different sample sizes using maximum coalition size for score QoI
Figure 7. Agreement of ShaRP for CSRankings when varying the sample size.
(a) Agreement of ShaRP across different coalitions sizes using maximum sample size for rank QoI
(b) Agreement of ShaRP across different coalitions sizes using maximum sample size for score QoI
(c) Agreement of ShaRP across different coalitions sizes using maximum sample size for pairwise QoI
Figure 8. Agreement of ShaRP for CSRankings when varying the coalition size.
7.2.2. Running time and quality of approximation
As discussed in Section 5.1, we implement two methods for approximating feature importance to reduce the running time: reduce the number of samples ($m$ in Algorithm 1) and restrict maximum coalition size ($|\mathcal{S}|$ in Algorithm 1). In addition to measuring the running time, we also measure and report explanation fidelity (Eq. 6 in Section 6.1) and method agreement (Eq. 9) to quantify the quality of the approximation. Figures 56,  7, and 8 present these results for the THE dataset, where exact computation was slowest and so there is most substantial room for improvement, as well as for the other two real datasets, CSR and ATP.
Figure 5a shows the speed-up when fewer samples $m$ are used for the computation. We observe that reducing the number of samples from 1378 (exact) down to 20, while keeping the maximum number of coalitions at 4 (largest possible for 5 features), speeds up rank QoI by a factor of 32, bringing the running time from 0.423 sec down to 0.013 sec. We see similar results for CSR 5b and ATP 5c, where speed-up is smaller because of the total number of features, but it is still significant, 6 for CSR (for 20 points down from 189) and 3 for ATP (for 20 points down from 86). Importantly, this performance gain does not come with a noticeable reduction in fidelity, which stays over 0.98 (out of 1) for all sample sizes in all experiments, Figure 6a6c.
Figures 5d5f and 6d6f show the speed-up and fidelity, respectively, when coalition size is bounded. We observe that speed-up is most substantial for all QoIs for coalitions of size 1, but that fidelity is lower for this approximation method. We observe fidelity values of at least 0.83 for rank, 0.91 for score, and 0.98 for pairwise. Fidelity further improves for rank and score for coalitions of size at most 2, reaching 0.87, 0.94, and 1, respectively. The difference between the approximation methods is expected because they affect different aspects of the sampling process. Sample size affects the number of samples we draw while maximum coalition size affects the number of features we can sample concurrently.
In Figures 7 and 8, we present the method agreement between the approximation and the exact computation for CSRankings (CSR). We omit method agreement results for the other datasets, where ShaRP  performs similarly. In 7a and 7b, we present agreement of the approximation when we vary sample size for the rank and the score QoI. We evaluate the agreement using the Jaccard Index (considering the top-2 features), Kendall’s tau distance, and the Euclidean distance of the feature vectors (converted to unit vectors). Here, we see that performance is similar for both QoIs. The Jaccard index is over 0.9 for any sample size, and is the distance metric with the worst performance for both QoIs. This is worth noting as shorter explanations are often considered more interpretable (Molnar, 2020). Agreement is similar or higher for all QoIs when we vary maximum coalition size, see Figure 8a8c.
In summary, reducing the number of samples and bounding coalition size are both effective at improving the running time, while retaining high explanation fidelity and method agreement in practice. Nonetheless, unless otherwise stated, we use exact computation in the remainder of this section to reduce variability in the reported results.
7.3. Qualitative Analysis
7.3.1. CS Ranking
Figure 9. Feature contributions to rank and score for the CSRankings dataset, aggregated over 10% strata. In this ranking, 189 computer science departments are ranked based on a normalized publication count of the faculty across 4 research areas: AI (green), Systems (orange), Theory (purple), and Interdisciplinary (pink). (a) Systems is the most important feature for an item’s rank in the top-20%, followed by AI. AI becomes more important for the rest of the ranking strata.  (b) Feature contributions to score are less informative than to rank: both capture the same relative feature importance for the top 20%; however, feature contributions become small and very similar as more items are tied for their score. (See rank vs. score plot on the top-right.)
In our next set of experiments, we generate and qualitatively analyze explanations for CS Rankings.
In Figure 9, we show feature contributions to rank and score QoI for the CS Rankings dataset, aggregated over 10% strata of the ranking. According to rank QoI in Figure 9a Systems is the most important feature in all strata, followed closely by AI. Both features have the most positive contributions in the top strata and the most negative in the bottom strata. Further, we observe that feature contributions to score are less informative than to rank: both capture the same relative feature importance for the top 20%; however, feature contributions become small and very similar as more items are tied or nearly-tied for their scores, making it nearly impossible to quantify feature importance across strata.
Figure 3, previewed in the introduction, presents the top-$k$ QoI for this dataset. The feature importance it shows is consistent with Figure 9b (score QoI), but the high positive impact of Systems on a department’s presence at the top-$k$ is even stronger pronounced. Additionally, unlike the score QoI, this top-$k$ QoI shows that Theory is impactful for getting into the top-$k$. Further, both top-$k$ and rank QoIs show that Systems and AI swap places in terms of importance, while score QoI does not detect this.
In Figure 3b3d we continue our analysis of the top-$k$ and consider the relative ranking of three universities: Georgia Tech in rank 5, Stanford in rank 6, and UMich in rank 7. We wish to understand why Georgia Tech is ranked higher than Stanford (Figure 3c), and why Stanford is ranked higher than UMich (Figure 3d). In both cases, Georgia Tech and UMich have lower values for all features except Systems. The Systems value of Georgia Tech is high enough to overcome the contributions of other features and rank it higher than Stanford. However, for UMich, we see that, while Systems is the most important feature in the top-10% stratum, it is not important enough to move UMich above Stanford. Note that adding the contributions for each feature adds up to the total difference in rank between each pair.
7.3.2. Score-based Ranking with Synthetic Data
Fixed scoring function, varying data distribution.
In this experiment, we illustrate that feature importance is impacted by the data distribution of the scoring features to a much greater extent than by the feature weights in the scoring function. Further, we show that feature importance varies by rank stratum. In Figure 10, we show rank QoI for 4 synthetic datasets with the same scoring function $f_{2}$.
We observe that, while the features have equal scoring function weights, their contributions to rank QoI differ for most datasets. In $D_{1}$, the Bernoulli-distributed $x_{2}$ determines whether the item is in the top or the bottom half of the ranking, while the Gaussian-distributed $x_{1}$ is responsible for the ranking inside each half. For $D_{2}$, the uniform $x_{1}$ has higher importance because it often takes on larger values than the Gaussian $x_{2}$. In $D_{4}$, $x_{1}$ and $x_{2}$ are negatively correlated, so when one contributes positively, the other contributes negatively. Only for $D_{3}$, with two uniform identically distributed features, the median contributions of both features are approximately the same within each stratum.
Additionally, we see that feature contributions differ per rank stratum. For example, for $D_{3}$, the medians show a downward trajectory across strata. This is because they quantify the expected change (positive or negative) in the number of rank positions to which the current feature values contribute. Also for $D_{3}$, feature contributions have higher variance in the middle of the range, because a 40-60% rank corresponds to many feature value combinations.
Figure 10. The rank QoI using $f_{2}=0.5\times x_{1}+0.5\times x_{2}$ for four datasets; $D_{1}$: $x_{1}\sim N(0.5,0.1)$, $x_{2}\sim Bern(0.5)$; $D_{2}$: $x_{1}\sim[0,1]$, $x_{2}\sim N(0.5,0.1)$; $D_{3}$: $x_{1}\sim[0,1]$, $x_{2}\sim[0,1]$; $D_{4}$: $x_{1}\sim N(0.5,0.05)$, $x_{2}\sim N(0.75,0.016)$, with -0.8 correlation. Feature contributions are different per rank stratum and data distribution.
Figure 11. Rank QoI for $D_{5}$: $x_{1}\sim N(0.5,0.1)$, $x_{2}\sim N(0.5,0.05)$. Subplots correspond to different scoring functions: $f_{1}=0.8\times x_{1}+0.2\times x_{2}$ (top), $f_{2}=0.5\times x_{1}+0.5\times x_{2}$ (middle), $f_{3}=0.2\times x_{1}+0.8\times x_{2}$ (bottom).
Fixed data distribution, varying scoring function.
In this experiment, we investigate the impact of the scoring function on rank and top-$k$ QoI for two datasets. In Figure 11, we use $D_{3}$ and see that the contributions to rank QoI vary depending on the scoring function. For $f_{1}$, $x_{1}$ is the only important feature (although it carries 0.8 — and not 1.0 — of the weight). This can be explained by the compounding effect of the higher scoring function weight and higher variance of the distribution from which $x_{1}$ is drawn. Between $f_{2}$ and $f_{3}$, features $x_{1}$ and $x_{2}$ switch positions in terms of importance, and show a similar trend, despite being associated with different scoring function weights (0.5 & 0.5 vs. 0.2 & 0.8). This, again, can be explained by the higher variance of $x_{1}$, hence, $x_{2}$ needs a higher scoring function weight to compensate for lower variance and achieve similar importance.
Figure 12. Top-$k$ QoI for $k=10\%$, $D_{2}$: $x_{1}\sim[0,1]$, $x_{2}\sim N(0.5,0.1)$, and $D_{3}$: $x_{1}\sim[0,1]$, $x_{2}\sim[0,1]$. Subplots correspond to different scoring functions: $f_{1}=0.8\times x_{1}+0.2\times x_{2}$ (top), $f_{2}=0.5\times x_{1}+0.5\times x_{2}$ (bottom).
Top-$k$ access.
Access to the top-$k$ is determined by the interaction between the scoring feature weights and the distributions of these features. The top-k QoI tells us how important each feature is when we consider only access to the top-$k$. A positive feature contribution signifies that changing the feature’s value will result in decreased chances of getting to the top-$k$. A very high (or very low) value shows that the changes are significant. Figure 12 illustrates this for datasets $D_{2}$ and $D_{3}$. When we consider two identical uniform features that have equal weights ($D_{2}$ under $f_{2}$), we first notice that their control of top-$k$ access is identical as expected. Additionally, we see that for the top-10, changing either feature would reduce access to the top-k (the values are both very positive). However, for each stratum up to the top-70%, changing either feature can contribute either positively or negatively.
When we consider two identical uniform features ($D_{1}$) one of which has a higher weight ($f_{1}$) or dataset $D_{2}$ (under either $f_{1}$ or $f_{2}$), we see again that for the top-10, changing either feature would reduce access to the top-k. Also, we see that how high the contributions are depends on the distributions. However, we see that for the top 10%-20% changing the second, less important feature would increase the chances of getting into the top-$k$. For the rest of the strata, with some variations depending on the dataset and function, changing the most important feature provides non-zero probability of moving to the top-$k$ and interestingly, this persists even for the lower strata. Evidence that items from lower strata can move to the top-$k$ under some scoring functions and feature distributions counters the assumption of Anahideh and Mohabbati-Kalejahi (2022) that changes in rank are localized.
7.4. Comparison to Other Methods
In this section, we compare explainability methods using the metrics described in Section 6. We restrict our analysis to a subset of the methods described in Section 2. The focus of ShaRP and the strength of Shapley values is individual explanations, for this reason, we do not compare to the methods that focus on providing global explanations. This excludes Yang et al., and Gale and Marian (Yang et al., 2018; Gale and Marian, 2020). To compare to HIL (Yuan and Dasgupta, 2023), we adapt their code to work with any score-based ranker and real data. We do not compare to PrefShap (Hu et al., 2022) which works only for pairwise datasets that use a generalized preferential kernel function. We compare to HRE (Anahideh and Mohabbati-Kalejahi, 2022) but use only four of their internal methods as provided by their public code base. Finally, we compare to SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b), due to their wide use and availability, even though they are not designed for ranking.
Implementation of HIL
HIL  (Yuan and Dasgupta, 2023) is the only other method that recommends the usage of ranks as a profit function for individual explanations in ranking. While this method is not general, we are interested in comparing it with our rank QoI. This was not straightforward because the method is available as a web app that works only for linear weight scoring functions and datasets of two Gaussian features. To compare the rank-relevance contributions introduced in that paper to the rank QoI, we adapted their method using their definitions and code. This implementation is available alongside our own. Further, we extended their method to work with the specific non-linear scoring function used by CS Ranking, by changing the way that Std rank and Std score (discussed below) are computed.
More specifically, because HIL  (Yuan and Dasgupta, 2023) works only with linear weight scoring functions, they do not provide a full Shapley values implementation but use the linear weights to approximate Shapley values assuming feature independence, see Corollary 1 in (Lundberg and Lee, 2017) and also (Strumbelj and Kononenko, 2014). This is a well-established method to compute Shapley values for linear weights also implemented by SHAP, so we do not compare with this part of the method. In addition, HIL defines two methods to acquire feature contributions: “standardized Shapley values” and “rank relevance Shapley values,” which we will call Std score and Std rank, respectively. Those are not calculated using the linear weight method described above, but rather directly from the weights, and without using the mean score or rank. For an item v, each feature $i$ contribution for Std score is $\phi_{i}=\frac{\beta_{i}\textbf{v}_{i}}{\sum_{\textbf{u}\in\mathcal{D}}f(u)}$, where $\beta_{i}$ is the weight for feature $i$. In other words, the contribution of each feature for each item is the score contribution of this feature over the sum of all scores for all items. Similarly, for Std rank, the contribution of feature $i$ for an item v is $\phi_{i}=\beta_{i}\textbf{v}_{i}\alpha_{\textbf{v}}$, where $\alpha_{\textbf{v}}$ is a scaling factor used to transform the score of the specific item to the rank of the specific item calculated as $\alpha_{\textbf{v}}=\frac{(\max_{r\in r_{\mathcal{D}}}(r)-r^{-1}_{\mathcal{D}(\textbf{v})})\sum_{\textbf{u}\in\mathcal{D}}f(u)}{\max_{r\in r_{\mathcal{D}}}(r)f(\textbf{v})}$. Note that neither of the two formulas is computing Shapley values; rather, they assign a contribution to the features based on the linear weights, and the score and rank. This implies that our rank QoI is the only rank QoI for Shapley values.
Details of HRE
For the purposes of comparison, we describe HRE (Anahideh and Mohabbati-Kalejahi, 2022) in more detail here. This method explains the ranked outcome of an item by fitting interpretable models in the immediate neighborhood of the item in the ranking (i.e., items that are ranked 5-10 consecutive positions above and below the item being explained). The interpretable models they make available through their public implementation are Decision Trees (DT), Linear Regression (LR), Ordinary Least Squares (OLS), and Partial Least Squares (PLS). Note that the feature-based explanations provided by HRE cannot be used to reconstruct the ranked outcome, and so we cannot compute fidelity for this method.
7.4.1. Fidelity
Table 3 presents the fidelity of all methods for which explanations can be used to reconstruct the ranked outcome, averaged across all items in all datasets. All methods are executed using their recommended settings to compute explanations for score QoI. Additionally, we include fidelity for ShaRP  for the rank QoI. Recall that ShaRP  is the only method that can compute an explanation for this QoI. We observe that LIME, SHAP and ShaRP  are all achieving high explanation fidelity, on average. HIL has reasonable fidelity for CSR but does not perform consistently on other datasets.
LIME SHAP ShaRP HIL
dataset score score score rank score
ATP 0.98 1.00 1.00 1.00 0.14
CSR 0.95 0.99 1.00 1.00 0.85
THE 0.94 0.97 1.00 1.00 0.64
Syn 0 0.95 0.99 1.00 1.00 0.37
Syn 1 0.95 0.99 1.00 1.00 0.29
Syn 2 0.95 0.99 1.00 1.00 0.35
Table 3. Fidelity across all methods across all datasets.
7.4.2. Agreement Between Explanations
Figure 13 presents the agreement between ShaRP  and all other methods for the rank and score QoIs in Figure 13a and Figure 13b, and the agreement between all methods averaged across all datasets in Figure 13c. We chose the rank and score QoIs for this comparison because we want to compare across all items of the dataset. and the methods we compare ShaRP to calculate neither pairwise nor presence in the top-k. Additionally, the results presented here are for Kendall’s tau distance between the explanations.
From these figures, we see that the explanations we will produce highly depend on the explanation method, regardless of the QoI. We see that ShaRP agrees the most with LIME and SHAP across both rank and score QoIs. HRE is expected to differ perhaps as it uses only localized information to produce explanations. However, when we are examining the similarity of explanations between HRE variants, we can still see that explanations are different for different HRE variants.
(a) Method Agreement between Sharp using the rank QoI and all methods across all datasets
(b) Method Agreement between Sharp using the score QoI and all methods across all datasets
(c) Method Agreement between all methods averaged across all datasets
Figure 13. Method Agreement
7.4.3. Sensitivity
(a) HRE DT
(b) HRE LR
(c) HRE OLS
(d) HRE PLS
(e) LIME
(f) SHAP
(g) ShaRP Score
(h) ShaRP Rank
(i) HIL Std Score
(j) HIL Std Rank
Figure 14. Sensitivity analysis for the CSRankings dataset for all methods. Each dot is a neighbor of the reference item. The x-axis shows the Euclidean explanation distance between the item and each neighbor, the y-axis the absolute rank difference between the reference item and each neighbor, and the hue shows how similar the features of each neighbor are to the reference item. ShaRP is the only method that can successfully provide similar explanations for items that are similar in feature space, and that also have similar outcomes.
Figure 14 presents the results of sensitivity analysis for all methods. In this figure, we evaluate the similarity of explanations for pairs of similar items. For each pair of items, we compute three distances: (1) Euclidean distance between the explanations (x-axis); (2) distance between the ranks of the two items (y-axis); and (3) Euclidean distance between the items in terms of their feature values (hue, where lighter means closer). To make the plot, we place one item (the reference item) at position (0,0) and use a scatter point for each other item (neighbor), indicating the distance in ranks and the distance of the explanations. The color of the scatter point indicates the distance between the features of the reference point and the neighbor. We then overlay the plots for all items in the dataset, so that all items are used as reference points.
Intuitively, we would expect items that are both similar in terms of their features and ranked near each other to have similar explanations. In other words, we would expect all points to be on or near the diagonal line $y=x$, with the hue getting darker as we move away from the origin. However, this is rarely the case in practice. For example, if one feature is far more important than the rest, then proximity in the Euclidean space of features will not be strongly correlated with explanation similarity.
Another example where this intuition does not hold is when two items that look very different in the feature space receive very similar outcomes. In both cases, we would like the explanations to be able to show this. Indeed, we should desire different explanations for items ranked close by, when their features are different. And we never want any explanations of items that are far from each other in the ranking to be very similar. This means that we want the plot to fill the space under the $y=x$ line and have a hue that gets darker as we move further from the origin.
Consider Figure 14, and observe that both ShaRP  with rank QoI (Figure 14h) and HIL-Std-Rank (Figure 14j, using our expanded definition of their method) are indeed able to produce similar explanations for similarly ranked items that have similar features, conforming to the shape we described above.  ShaRP  is performing slightly better in clustering the points into denser similarity groups. On the other hand, SHAP (Figure 14f), a score-based method, bases its explanations mostly on the feature distances. Indeed, the plot consists of clear bands that get darker away from 0 and below the diagonal. Nevertheless, items that are ranked very differently but have similar features receive almost identical explanations. ShaRP  with score QoI in Figure 14g is almost identical to SHAP, as we also saw in method agreement, and HIL using the expanded score definition in Figure 14i also performs similarly. LIME (Figure 14e), which is also score-based, is producing a plot that suggests that the method uses both feature distance and ranking distance, but is still unable to differentiate between the explanations as clearly as the rank-based methods. Finally, the explanation distance of all HRE methods does not appear to depend on the ranking distance or the feature distance as we see items with similar and dissimilar explanations across all ranks and hues. This is expected, as the explanations depend only on close neighbors (that can have different features) and there is no guarantee that a different cluster of explanations for a further neighborhood will be different from the neighborhood of the reference point.
To further compare ShaRP  with rank QoI and HIL with Std rank, we present Figure 15. Even though both methods are appropriate for the ranking task we are examining, in this figure, we see that ShaRP  with rank QoI Figure 15a,  15c, and  15e can capture the full range of different ranks and features, and that groups the items more successfully. HIL with Std rank cannot capture the difference of feature values for ATP (Figure 15b), or the similarly ranked items that have different feature values for the Synthetic experiment (the middle area close to the $x$-axis of Figure 15f. Both methods perform similarly for THE (Figures 15c and  15d).
In summary, explanations for the rank QoI, which we are introducing in this paper, are able to more accurately explain ranking tasks compared to other local feature-based explanation methods.
(a) ShaRP Rank ATP
(b) HIL Std Rank ATP
(c) ShaRP Rank THE
(d) HIL Std Rank THE
(e) ShaRP Rank Synthetic-0
(f) HIL Std Rank Synthetic-0
Figure 15. Sensitivity analysis for the ATP, THE, and Synthetic dataset 0 for the methods using the rank QoI.
8. Conclusions, Limitations, and Future Work
We presented ShaRP , a comprehensive framework for quantifying feature importance for selection and ranking. Because of the impact rankers have on individuals, organizations, and population groups, there is a need to understand them: to know whether the decisions are abiding by the law (auditing and compliance tasks), to help individuals improve their ranked outcomes (recourse tasks), and to design better ranking procedures (design task). Our work is motivated by the foundational interpretability needs of these tasks.
We showcased the usefulness of ShaRP  for qualitative analysis of an impactful real dataset and task — the ranking of Computer Science departments. We augmented this analysis with a thorough evaluation of our methods on numerous real and synthetic datasets, showing that the quantities of interest we define bring interesting — and complementary — insights about the data that cannot be captured simply by the relationship between the scores and the ranks. We showed that feature importance depends on the properties of the data distribution even if the scoring function remains fixed, and, further, that feature importance exhibits locality. Finally, we compared our method to other local feature-based explanation methods and showed that it compares to them favorably.
Limitations and future work.
A limitation of our work is that our evaluation of the usefulness of explainability did not include a user study. Understanding how humans make sense of feature importance in ranking and act on their understanding, is in our immediate plans.
We also plan to support additional quantities of interest, incorporate learning-to-rank methods into our evaluation, and develop a comprehensive benchmark that compares both local and global feature importance methods for ranking in terms of their usability, expressiveness, and performance.
Acknowledgements.
We would like to thank Tilun Wang for her contributions to the code base that supported an earlier version of this draft. We would also like to thank Lucius Bynum and Falaah Arif Khan for their input and advice while developing this work. This research was supported in part by NSF Awards No. 2312930 and 2326193, and by the NYU Tandon Center for Responsible AI, through their Responsible AI for Ukraine program, r-ai.co/ukraine.