### v4

# ShaRP: Explaining Rankings and Preferences with Shapley Values

Abstract.
Algorithmic decisions in critical domains such as hiring, college admissions, and lending are often based on rankings. Given the impact of these decisions on individuals, organizations, and population groups, it is essential to understand them—to help individuals improve their ranking position, design better ranking procedures, and ensure legal compliance. In this paper, we argue that explainability methods for classification and regression, such as SHAP, are insufficient for ranking tasks, and present ShaRP—Shapley Values for Rankings and Preferences—a framework that explains the contributions of features to various aspects of a ranked outcome.
ShaRP computes feature contributions for various ranking-specific profit functions, such as rank and top-$k$, and also includes a novel Shapley value-based method for explaining pairwise preference outcomes. We provide a flexible implementation of ShaRP, capable of efficiently and comprehensively explaining ranked and pairwise outcomes over tabular data, in score-based ranking and learning-to-rank tasks. Finally, to evaluate ShaRP and compare it with other explainability methods, we define ranking-specific explanation metrics and conduct an extensive experimental analysis, demonstrating the framework’s flexibility and efficiency.
ranking, interpretability, feature importance, Shapley values, evaluation, responsible data management, responsible AI
1. Introduction
Rankings produced by data-driven algorithmic systems now influence a myriad of socio-technical applications, as part of automated or semi-automated decision-making, and with direct consequences to people’s lives and aspirations. An algorithmic ranker, or a ranker for short, takes a database of candidates as input and produces a permutation of these candidates as output, see Figure 1 for an example. We refer to the output of a ranker as a ranked outcome or simply a ranking. As an alternative to the full permutation, the best-ranked $k$ candidates, or the top-$k$, may be returned in rank order or as a set. In the latter case, we are dealing with a selection task, which is a special case of ranking.
Algorithmic rankers are broadly used to support decision-making in critical domains, including hiring and employment, school and college admissions, credit and lending, and, of course, college ranking. Because of the impact rankers have on individuals, organizations, and population groups, there is a need to understand them: to know whether the decisions are correct and legally compliant (auditing tasks), to help individuals improve their ranked outcomes (recourse tasks), and to design better ranking procedures (design tasks). To make progress towards these tasks, we need ways to explain and interpret ranked outcomes. In this paper, we present ShaRP  —Shapley for Rankings and Preferences—a framework that explains the contributions of features to different aspects of a ranked outcome, and that can support all these critically important tasks.
There are two types of rankers: score-based and learned. In score-based ranking, a given set of candidates is sorted on a score, which is typically computed using a simple formula, such as a weighted sum of attribute values (Zehlike et al., 2022a). In supervised learning-to-rank (LtR), a preference-enriched set of candidates is used to train a model that predicts rankings of unseen candidates (Li, 2014). Score-based rankers are a prominent example of the so-called “interpretable models” (Rudin, 2019): the scoring function, such as $Y_{1}=0.4\times gpa+0.4\times sat+0.2\times essay$ in a college admissions scenario, is based on a (normative) a priori understanding of what makes for a good candidate. And yet, despite being syntactically “interpretable,” a score-based ranker may not be “explainable,” in that, the designer of the ranker or the decision-maker who uses it may be unable to accurately predict or understand its output ((Miller, 2019; Molnar, 2020)). We illustrate this with an example.
name gpa sat essay $f$ $g$
Bob 4 5 5 4.6 5
Cal 4 5 5 4.6 5
Dia 5 4 4 4.4 4
Eli 4 5 3 4.2 3
Fay 5 4 3 4.2 3
Kat 5 4 2 4.0 2
Leo 4 4 3 3.8 3
Osi 3 3 3 3.0 3
$r_{\mathcal{D},f}$
Bob
Cal
Dia
Eli
Fay
Kat
Leo
Osi
$r_{\mathcal{D},g}$
Bob
Cal
Dia
Eli
Fay
Leo
Osi
Kat
Figure 1. (a) Dataset $\mathcal{D}$ of college applicants, scored on $gpa$, $sat$, and $essay$. (b) Ranking $r_{\mathcal{D},f}$ of $\mathcal{D}$ on $f=0.4\times gpa+0.4\times sat+0.2\times essay$; the highlighted top-4 candidates will be interviewed and potentially admitted. (c) Ranking $r_{\mathcal{D},g}$ on $g=1.0\times essay$; the top-$4$ coincides with that of $r_{\mathcal{D},f}$, signifying that $essay$ has the highest importance for $f$, despite carrying the lowest weight in the scoring function.
Example 0.
Consider a dataset $\mathcal{D}$ of college applicants in Figure 1, with scoring features $gpa$, $sat$, and $essay$. Very different scoring functions $f=0.4\times gpa+0.4\times sat+0.2\times essay$ and $g=1.0\times essay$ induce very similar rankings $r_{\mathcal{D},f}$ and $r_{\mathcal{D},g}$, with the same top-$4$ items appearing in the same order, apparently because $essay$ is the feature that is best able to discriminate between the top-$4$ and the rest, and that determines the relative order among the top-$4$.
This example illustrates that a scoring feature’s weight may not reflect its impact on the ranked outcome. This occurs because rankings are inherently relative, while feature values and computed scores are absolute—knowing an item’s score reveals little about its position relative to others. The lack of independence between per-item outcomes renders existing feature importance methods for classification and regression (Covert et al., 2020; Datta et al., 2016; Guidotti et al., 2019; Lundberg and Lee, 2017; Mohler et al., 2020; Ribeiro et al., 2016b; Strumbelj and Kononenko, 2010) insufficient for ranking tasks. These methods assess how a feature influences an item’s score, but a feature may significantly affect an item’s score while leaving its rank unchanged. Consider the following example.
Example 0.
Consider Figure 1 and suppose that Dia’s essay score increases from 4 to 5, thus increasing the scores computed with both $f$ (4.4 to 4.6) and $g$ (4 to 5). However, Dia’s rank remains unchanged.
Changes in score do not necessarily lead to changes in rank because, in selection and ranking, an item’s outcome v depends on the outcomes of other items in $\mathcal{D}\setminus\{\textbf{v}\}$. For example, only one item can occupy a given rank, and exactly $k$ items can appear at the top-$k$. Thus, any explainability method that measures score changes can only partially explain rank changes. This highlights that interpretability for ranking tasks requires measuring the features’ impact on quantities beyond the score, such as rank or top-$k$ presence. We preview these results for CS Rankings in Figure 2, where feature importance for score in 2a and rank in 2b yield markedly different explanations. We discuss these findings in detail in Section 4.
(a) Score QoI: Interdisciplinary is the most important feature, followed by AI and Systems, while Theory negatively impacts the score.
(b) Rank QoI: Systems is the most important feature, followed by AI and Interdisciplinary. Theory is minimally but positively impacting the rank.
Figure 2. Feature importance for Texas A&M in CS Rankings.
In summary,
ranking fundamentally differs from classification and regression, as reflected in the extensive literature on learning-to-rank (Li, 2014). Fairness in ranking has also been addressed separately from fairness in classification (Zehlike et al., 2022a, b). Likewise, interpretability methods must be tailored for ranking, as we do in this work. The key insight is that the impact of scoring features on ranking must be measured in a way that accounts for the lack of independence between per-item outcomes. This work formalizes and builds upon this intuition, making four contributions.
Our first contribution is the formalization of several natural profit functions for computing Shapley values in ranking, capturing feature contributions to an item’s score, rank, or top-$k$ presence
We build on the Quantitative Input Influence (QII) framework by Datta et al. (2016), which quantifies feature importance in classification using Shapley values (Shapley et al., 1953), a standard method for revenue division in cooperative games. In QII, an item’s features (players) influence its classification outcome (profit function) within a cooperative game. We adopt QII as a flexible foundation for exploring different profit functions in Shapley value explanations. QII refers to profit functions as Quantities of Interest (QoIs), a term we use in this work.
Our second contribution is a new Shapley value method for explaining pairwise outcomes (preference pairs).
A class of Shapley value methods generates explanations by comparing an item to a “baseline” item considered “neutral” (Lundberg and Lee, 2017; Chen et al., 2023). Instead of using a fixed baseline item, we dynamically adjust it for each pair $u\prec v$, computing feature importance for $v$ using $u$ and vice versa. This approach yields feature contributions that naturally explain the difference between the compared items’ outcomes.
Our third contribution is a robust and extensible open-source library called ShaRP , the only available library that uses Shapley values for explaining ranked outcomes over tabular data.
Our implementation includes both exact and approximate QoI computation to achieve interactive performance while retaining explanation quality. As part of ShaRP , we are also making an implementation of the original QII framework available to the community.
Our fourth contribution is a methodology for evaluating ranking explainability methods for tabular data, supported by extensive experiments.
We define evaluation metrics and conduct qualitative analyses on real and synthetic datasets, showing that feature importance for all QoIs depends on data distribution properties and ranking position, in addition to the scoring function. Additionally, we demonstrate that the rank-aware QoIs we introduce provide unique and complementary insights beyond score-rank relationships. We also present a large-scale quantitative evaluation of ShaRP , highlighting the scalability of approximations and comparing it to other methods.
2. Related Work
Local feature-based explanations
Ribeiro et al. (2016b) introduced LIME, which explains complex classification models locally using interpretable models like sparse linear regressions. Lundberg and Lee (2017) proposed SHAP, which leverages Shapley values to explain various classification and regression models. Both methods are extensively implemented in software libraries and can be used to explain an item’s score (score QoI in our terminology).
Feature-based explanations for ranking
Yang et al. (2018) introduced a “nutritional label” for score-based rankers, featuring two visual widgets: “Recipe,” which displays scoring feature weights, and “Ingredients,” which highlights features (scoring or not) with the strongest Spearman’s rank correlation with the score. They observed that a feature’s weight in “Recipe” often does not correspond to its correlation with the score. This approach provides global feature-based explanations, identifying which features influence the ranked outcome for all items. In contrast, we focus on local explanations, highlighting feature importance for a specific item or pair of items.
Gale and Marian (2020) introduced “participation metrics” for score-based rankers, measuring how scoring features contribute to an item’s presence in the top-$k$. Their key metric, “weighted participation,” attributes an item’s inclusion in the top-$k$ to its feature values, scoring function weights, distance from maximum possible values, and whether a feature value exceeds the $k^{th}$ highest value. This quantity is aggregated across all top-$k$ items, producing a single participation value per feature. The main distinction between our method and Gale and Marian (2020) is that we provide local (per-item) explanations, whereas their approach offers group-wise explanations for the top-$k$. Our top-$k$ QoI is similar to weighted participation but focuses on local explanations, which can optionally be aggregated across the top-$k$.
Yuan and Dasgupta (2023) developed an interface for sensitivity analysis of ranked synthetic datasets with two normally distributed features and a linear scoring function. They quantify feature importance by subtracting a feature’s value from its mean, which, under their assumptions, approximates Shapley values (Lundberg and Lee, 2017). They also propose two alternative methods for quantifying feature contributions using the linear weights and the item’s score/rank. To compare with their approach, we re-implemented and significantly extended their method to support more than two features, arbitrary feature distributions, and flexible scoring functions.
Anahideh and Mohabbati-Kalejahi (2022) provide local feature-based explanations using multiple methods, including Shapley values computed with SHAP (Lundberg and Lee, 2017), and apply these methods locally to items ranked around the one being explained. They justify this approach by noting that in repeating competitions, such as school rankings, items tend to maintain similar ranks. Like Anahideh and Mohabbati-Kalejahi (2022), we find that feature importance varies significantly across rank strata. However, contrary to their assumptions, we show that slight changes in feature values can cause substantial rank shifts, extending far beyond the local neighborhood.
Moskovitch et al. (2023) developed DEXER to detect groups that may be under- or overrepresented in the top-$k$ compared to their proportion in the input. To analyze the causes of these disparities, they fit a linear regression model to ranks (treating them as scores) and compute feature importance using SHAP (Lundberg and Lee, 2017). In contrast, we fully adapt Shapley values for rank-specific outcomes in our ShaRP  framework. We compare the explanations generated by ShaRP  and DEXER in Section 8.2.
Similarly, Pastor et al. (2021) proposed a method that uses itemsets to automatically detect under- or overrepresented subgroups, where each subgroup is defined by a subset of the attributes. Like our approach, they define a ranking profit function, but they use it to quantify each attribute’s contribution to the itemset’s outcomes.
Finally, Hu et al. (2022) introduced PrefSHAP, a method for explaining pairwise preferences of dueling data modeled using the generalized preferential kernel (GPM) from their prior work (Chau et al., 2022). GPM is ideal for modeling inconsistent preferences that do not lead to a global ranking. Similar to our paper, they show that score-based Shapley values are insufficient to explain GPM preferences, and they define a new Shapley value profit function for this type of data. To compute Shapley value contributions, they apply the transformations for every coalition on both items at the same time and evaluate how the preference changes. Like Hu et al. (2022), we argue that ranking-specific quantities of interest (QoIs) are essential for explaining rankings and preferences, and introduce a broader set, including score, rank, and top-$k$, and a new pairwise explainability method. We do not compare experimentally with PrefSHAP because our focus is on preferences induced by score-based rankers or LtR, whereas PrefSHAP applies only to a specific class of kernel models.
Shapley-based explanations in Information Retrieval (IR)
Concurrently with our work, Heuss et al. (2024) and Chowdhury et al. (2024) published preprints adapting Shapley values for explaining ranked outcomes in IR. Both methods generate feature importance for the entire ranking rather than individual ranked items, therefore producing global Shapley-based explanations. To achieve this, they use SHAP (Lundberg and Lee, 2016), varying the same subset of features in each coalition for all items (documents). In contrast, our explanations are local (explaining outcomes for one item at a time), and so our work is not directly comparable to (Heuss et al., 2024; Chowdhury et al., 2024). Furthermore, our focus is on explaining rankings over tabular data, rather than over document collections. Other recent work in IR explored the use of LIME to explain ranked outcomes (Verma and Ganguly, 2019; Singh and Anand, 2018; Chowdhury et al., 2023), and introduced baseline document construction techniques for DeepSHAP (Lundberg and Lee, 2017) to improve explanation quality (Singh and Anand, 6 29).
In summary,
we share motivation with all these lines of work but take a leap, describing the first comprehensive Shapley-value-based framework for explaining ranking and pairwise preference outcomes.
3. Preliminaries and Notation
Ranking.
Let $\mathcal{A}$ denote an ordered collection of features (equiv. attributes), and let $\mathcal{D}_{\mathit{}}$ denote a set of items (equiv. points or candidates). An item $\textbf{v}=(v_{1},\dots,v_{d})\in\mathbb{R}^{d}$ assigns values to $|\mathcal{A}|=d$ features, and may additionally be associated with a score. Score-based rankers use a scoring function $f(\textbf{v})$ to compute the score of v. For example, using $f_{1}(\textbf{v})=0.4\times gpa+0.4\times sat+0.2\times essay$, we compute $f(\texttt{Bob})=4.6$ and $f(\texttt{Leo})=3.8$.
A ranking $r_{\mathcal{D}}$ is a permutation over the items in $\mathcal{D}_{\mathit{}}$. Letting $n=\left|\mathcal{D}_{\mathit{}}\right|$, we denote by $r_{\mathcal{D}}=\langle\textbf{v}_{1},\ldots,\textbf{v}_{n}\rangle$ a ranking that places item $\textbf{v}_{i}$ at rank $i$. We denote by $r_{\mathcal{D}}(i)$ the item at rank $i$, and by $r_{\mathcal{D}}^{-1}(\textbf{v})$ the rank of item v in $r_{\mathcal{D}}$. In score-based ranking, we are interested in rankings induced by some scoring function $f$. We denote these rankings $r_{\mathcal{D},f}$. For example, in Figure 1b, $r_{\mathcal{D},f}(1)=\texttt{Bob}$, $r_{\mathcal{D},f}^{-1}(\texttt{Leo})=7$. We assume that $r_{\mathcal{D},f}^{-1}(\textbf{v}_{1})<r_{\mathcal{D},f}^{-1}(\textbf{v}_{2})<\dots<r_{\mathcal{D},f}^{-1}(\textbf{v}_{n})$, where smaller rank means better position in the ranking.
[…]
Our goal is to explain the importance of features $\mathcal{A}$ to the ranking $r_{\mathcal{D},f}$. We will do so using Shapley values (Shapley et al., 1953).
Shapley values
For a set $\mathcal{N}$ of $n$ players, and a value function $f$ that assigns a profit to any subset (or coalition) $\mathcal{S}$ of players, $f:2^{n}\rightarrow\mathbb{R}$, where $f(\emptyset)=0$, the Shapley value of player $i$ is:
(1) $\phi_{i}(f)=\sum_{\mathcal{S}}\frac{|\mathcal{S}|!(n-|\mathcal{S}|-1)!}{n!}(f(\mathcal{S}\cup\{i\})-f(\mathcal{S}))$
We will use Shapley values to explain ranked outcomes using the set of features $\mathcal{A}$ as the players, and the outcome (or the quantity of interest, QoI) as the payoff function. In addition to the definition of players and the payoff function, Shapley values require the quantification of the payoff over a subset of the players. This, in turn, requires some way to estimate the payoff over a subset of the features. Consequently, for any Shapley value implementation, a method of feature removal or masking is required (Covert et al., 2021; Chen et al., 2023).
A common method (e.g., used in SHAP (Lundberg and Lee, 2017; Covert et al., 2021)), for a coalition $\mathcal{S}\subseteq\mathcal{A}$, is to marginalize out the features not in the coalition $\mathcal{A}\setminus\mathcal{S}$ and draw values from the marginal distributions of the subset of features in $\mathcal{S}$ jointly, often referred to as the “marginal” approach. Another alternative (e.g., used in QII (Datta et al., 2016)) is to draw values of each feature in $\mathcal{S}$ independently from its marginal distribution, often referred to as the “product of marginals” approach. Another approach is called “baseline” and instead of sampling the features in the coalition, they are replaced with the feature values of a specific fixed sample (Lundberg and Lee, 2017). Here, we choose the marginal approach for our implementation and take inspiration from the baseline approach for our pairwise method. In Section 7, we show how both can be implemented using one algorithm.
Let $\textbf{v}_{\mathcal{S}}$ denote a projection of v onto $\mathcal{S}$. In the example in Figure 1, $(\texttt{Bob},\texttt{4},\texttt{5},\texttt{2})_{\{name,gpa\}}=(\texttt{Bob},\texttt{4})$. We draw values from the marginal distributions of the subset of features in $\mathcal{S}$ as follows. Let U $=\langle\textbf{u}_{1},\dots,\textbf{u}_{m}\rangle$ denote a random vector of $m$ items sampled from $\mathcal{D}$. For a subset of features $\mathcal{S}\in\mathcal{A}$, let $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}=\langle\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{1})_{\mathcal{S}},\dots,\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{m})_{\mathcal{S}}\rangle$ denote a vector of items, in which each $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{i})_{\mathcal{S}}$ takes on the values of the features in $\mathcal{S}$ from $\textbf{u}_{i}$, and the values of the remaining features $\mathcal{A}\setminus\mathcal{S}$ from v. We calculate Shapley values using this set of features $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$, note that if $m=|\mathcal{D}-1|$ we use the entire dataset $\mathcal{D}\setminus\textbf{v}$ to calculate the exact Shapley values.
Shapley values satisfy several natural axioms, including efficiency, symmetry, dummy, and additivity (Shapley et al., 1953), with additional useful properties, such as monotonicity, following from these axioms (Lundberg and Lee, 2017). Efficiency states that the sum of the contributions of all features for item v equals the difference between the outcome $f(\textbf{v})$ and the average outcome: $\sum_{i\in\mathcal{A}}\varphi_{i}(f,\textbf{v})=f(\textbf{v})-\mathbb{E}_{\textbf{X}}[f(\textbf{X})]$ (Covert et al., 2020; Molnar, 2020). Using this property, the explanation can be used to reconstruct the outcome. We will use the efficiency property to define the fidelity metric for comparing explanations (Section 6).
4. Quantities of Interest for Ranking
The first contribution of our work is that we define QoIs that are appropriate for ranked outcomes. In addition to the score, we introduce the rank and top-$k$ QoIs. We use the notation for the marginal feature removal approach in this section but note that the QoIs we introduce can be used with any feature removal approach.
Score QoI
The Shapley value function for the score QoI is:
(2) $QoI_{f,\textbf{v}}(S)=\mathop{\mathbb{E}}_{\textbf{U}_{\mathcal{S}}}[f(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
This QoI captures the impact of an item’s features on its score. This is the QoI used by the popular feature-based explanation methods such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016a). To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the score over $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$ using the random variable $\textbf{U}_{\mathcal{S}}$ that takes values from the marginal distributions of the set of features in $\mathcal{S}$.
Figure 3. Feature contributions to rank and score for the CSRankings dataset, aggregated over 10% strata. In this ranking, 189 computer science departments are ranked based on a normalized publication count of the faculty across 4 research areas: AI (green), Systems (orange), Theory (purple), and Interdisciplinary (pink). (a) Systems is the most important feature for an item’s rank in the top-20%, followed by AI. AI becomes more important for the rest of the ranking strata.  (b) Feature contributions to score are less informative than to rank: both capture the same relative feature importance for the top 20%; however, feature contributions become small and very similar as more items are tied for their score. (See rank vs. score plot on the top-right.)
Rank QoI
The Shapley value payoff function for the rank QoI is:
(3) $QoI_{f,\textbf{v},\mathcal{D}}(S)=\mathop{\mathbb{E}}_{\textbf{U}_{\mathcal{S}}}[r^{-1}_{\mathcal{D}^{\prime},f}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. This QoI evaluates the impact of an item’s features on its rank. To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the rank over $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$ using the random vector $\textbf{U}_{\mathcal{S}}$ using the random variable $\textbf{U}_{\mathcal{S}}$ that taakes values from the marginal distributions of the set of features in $\mathcal{S}$.
Top-$k$ QoI
The Shapley value payoff function to quantify the impact of an item’s features on its presence or absence among the top-$k$ is stated similarly as rank QoI:
(4) $QoI_{f,\textbf{v},\mathcal{D}}(S)=\mathop{\mathbb{E}}_{\textbf{U}_{\mathcal{S}}}[\mathbbm{1}_{r_{\mathcal{D}^{\prime},f}(1\ldots k)}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. The difference with rank QoI (Equation 3) is that here we compute the expectation over the indicator function that returns 1 if an item’s rank is at most $k$ and 0 otherwise. This QoI allows us to quantify how each feature contributed to getting the item into the top-$k$.
Shapley values for ranking
To compute Shapley values for the QoIs we defined, we need to apply Equation 1 on the QoIs. Following the QII notation, we define the iota function $\iota$ as the difference between the QoIs of $\mathcal{S}\cup i$ and $\mathcal{S}$, the coalition including feature $i$ and the coalition.
(5) $\iota_{f,\mathbf{v},\mathcal{D}}(i,\mathcal{S})=\alpha(QoI_{f,\mathbf{v},\mathcal{D}}(\mathcal{S}\cup i)-QoI_{f,\mathbf{v},\mathcal{D}}(\mathcal{S}))$
where, the QoI can be any of the QoIs defined in this section, and $\alpha\in\{-1,1\}$ is a multiplier that adjusts the order of QoI terms. This is necessary when consider QoIs such as rank, where smaller values are preferable. In this case, we set $\alpha=-1$ to adjust the $\iota$ function accordingly.
Using this notation, we can define Shapley values for ShaRP:
(6) $\phi_{i}(f,\mathbf{v},\mathcal{D})=\sum_{\mathcal{S}}\frac{|\mathcal{S}|!(n-|\mathcal{S}|-1)!}{n!}\iota_{f,\mathbf{v},\mathcal{D}}(i,\mathcal{S})$
Case Study: QoIs for CSRankings
We now review some of the local feature-based explanations of ranked outcomes generated with ShaRP  over CS Rankings, a real dataset that ranks 189 Computer Science departments in the US based on a normalized publication count of the faculty across 4 research areas: AI, Systems, Theory, and Interdisciplinary (Berger, 2023), see Appendix  A for a detailed description of the dataset and the ranker. Here, our goal is to highlight some interesting insights about the data that these explanations can surface and, further, to show that these insights can be very different depending on what kind of outcome is being explained.
In Figure 3, we show feature contributions to rank and score QoI for the CS Rankings dataset, aggregated over 10% strata of the ranking. According to rank QoI in Figure 3a Systems is the most important feature in all strata, followed closely by AI. Both features have the most positive contributions in the top strata and the most negative in the bottom strata. Further, we observe that feature contributions to score are less informative than to rank: both capture the same relative feature importance for the top 20%; however, feature contributions become small and very similar as more items are tied or nearly tied for their scores, making it nearly impossible to quantify feature importance across strata.
Similarly, Figure 4a presents the aggregated feature contributions to the top-$k$ in CS Rankings. In this figure, the contribution of each of the four features to the top-$k$ is summarized for the highest-ranking 10% of the departments, followed by the next 10%, and so on. We again observe that a strong publication record in systems is the most important feature for placing a CS department among the top-10%, significantly ahead of AI. The feature importance it shows is consistent with Figure 3b (score QoI), but the high positive impact of Systems on a department’s presence at the top-$k$ is even stronger pronounced. Additionally, unlike the score QoI, this top-$k$ QoI shows that Theory is impactful for getting into the top-$k$.
Figure 2, previewed in the Introduction, presents a feature-based explanation for Texas A&M in CS Rankings, where it has a score of 3.941 and is ranked in position 34. Figures 2a and 2b show feature importance for the score and rank QoIs, respectively. Both are presented as waterfall plots (Lundberg and Lee, 2017). This plot displays how each feature contributed to the difference of the item’s outcome $f(x)$ from the mean outcome $\mathbb{E}[f(X)]$. For example, Texas A&M’s 0.16 AI value helped it score 0.45 higher than the mean score of 2.722. Features in red in that figure positively impact the QoI and in blue negatively. The sign depends on whether a higher (score) or a lower (rank) QoI is better. The sum of feature importance values equals the difference between the item’s QoI and the mean. Finally, the magnitude of feature importance depends on the QoI and is hence not directly comparable across QoIs.
According to Figure 2a, “Interdisciplinary” is the most (positively) impactful feature on the score of Texas A&M, followed by “AI” and then by “Systems”, while “Theory” contributes negatively. In contrast, all features contribute positively to the rank of Texas A&M, with “Systems” being the most impactful. This observation illustrates that different QoI are useful for answering different questions. If the goal is to improve the score, then Texas A&M should invest in interdisciplinary and AI research. However, if the goal is to improve the rank, then it’s better to prioritize systems research. Why are the score and rank QoI explanations different? The reason is that a change in score may not result in a change in rank. Specifically, for a rank to change, the change in the score must be large enough to surpass the score of the item ranked directly above.
Another key aspect of these plots is the color of each feature, which indicates whether a feature contributes positively or negatively to the outcome. The contributions add up to the difference from the mean outcome so, their magnitude and color vary depending on the QoI. For example, the average score is influenced by outliers, while rank is not (by construction), so we expect the color of the contributions to be very different. Indeed, in CS Rankings, over 70% of departments have scores below the mean. As a result, when using the score QoI, many or all of their features appear to contribute negatively. This highlights that the meaning of positive and negative contributions is dependent on the chosen QoI. See Figure 9b in the Appendix for the score vs. rank distribution for this dataset.
(a) Feature contribution to the top-$k$ QoI, for $k=10\%$. Systems is the most important feature, followed by Interdisciplinary and AI.
Institution AI Systems Theory Inter. Rank
Georgia Tech 28.5 7.8 6.9 10.2 5
Stanford 36.7 5.4 13.3 11.5 6
UMich 30.4 9.0 9.3 5.9 7
(b) Feature values and rank of three highly ranked departments: Georgia Tech, Stanford, and UMich.
(c) Pairwise method using the rank QoI: Georgia Tech ranks higher than Stanford because of its relative strength in Systems.
(d) Pairwise method using the rank QoI: Stanford ranks higher than UMich despite Stanford’s relative weakness in Systems.
Figure 4. Feature importance for the top-$k$ QoI (i.e., selection) for CS Rankings in 4a, with further analysis of the relative orders among two pairs of departments in 4c and 4d.
5. Shapley value Method for Pairwise Explanations
[…]
In Section 3 we discussed three different feature removal techniques for Shapley values. Here, we introduce a new Shapley value explanation method that uses the baseline feature removal technique. Note that, as we explained in Section 4, all QoIs can be used with any feature removal method so, any QoI can be used to calculate the pairwise explanation.
Baseline feature removal techniques select one “neutral” item as the “baseline” and then compare all other items to it. The benefit of these methods is that the exact feature contributions can be computed without any sampling. The disadvantage is that it is often hard to select the baseline sample because different baseline samples create different feature attributions and, in most contexts, it is hard to identify a “neutral” or “average” item. As an example, in related work, we mentioned  (Singh and Anand, 6 29) that attempts to identify a good baseline input document for DeepSHAP in IR. As another example, the baseline implementation of SHAP (Lundberg and Lee, 2017) uses the all-zeroes item as the baseline sample. While selecting a baseline sample is not simple in most cases, we find that the baseline feature removal technique is a natural fit when we are explaining the difference in outcomes between two items v and u.
When explaining the pairwise outcome of two items v and u, we are going to generate an explanation for one item using the other as the baseline. In other words, for coalition $\mathcal{S}$, we will be creating the hybrid item $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}$. Note that we do not need the feature distributions or any other parameters for this method. Additionally, note that we are not selecting a fixed item as the baseline but we dynamically change it depending on the pair of items we want to compare. This definition has a natural interpretation, the feature importance of a pairwise explanation amounts to the difference between the outcome of the two items. According to the property of efficiency (see Section 3) we have: $\sum_{i\in\mathcal{A}}\varphi_{i}(f,\textbf{v})=f(\textbf{v})-\mathbb{E}_{\textbf{X}}[f(\textbf{X})]=f(\textbf{v})-\frac{1}{2}(f(\textbf{v})+f(\textbf{u}))=\frac{1}{2}(f(\textbf{v})-f(\textbf{u}))$.
The Shapley value of v in comparison to u is defined as:
(7) $\phi_{i}(f,\mathbf{v}\succ\textbf{u})=\sum_{\mathcal{S}}\frac{|\mathcal{S}|!(n-|\mathcal{S}|-1)!}{n!}\iota_{f,\mathbf{v},\textbf{u}}(i,\mathcal{S})$
Note that Eq. 7 differs from Eq. 6 in setting $\mathcal{D}=\{\textbf{u}\}$. Hence, in Section 7 we are able to implement both the baseline and the marginal feature removal techniques using one algorithm. Because pairwise preferences are of special interest to ranking tasks, we will only be using rank as the QoI for the pairwise method in the rest of the paper.
Case Study: Explanations of Pairwise Outcomes for CSRankings
In Figure 4b4d we continue our analysis of the top-$k$ and consider the relative ranking of three universities: Georgia Tech in rank 5, Stanford in rank 6, and UMich in rank 7. We use pairwise Shapley explanations that can clarify rank differences between two items. Consider Fig. 4c, where we explain the pairwise outcome for Georgia Tech $\textbf{v}_{\text{Georgia Tech}}=(28.5,7.8,6.9,10.2)$ vs $\textbf{u}_{\text{Stanford}}=(36.7,5.4,13.3,11.5)$. To explain the pairwise outcome difference, for each coalition, we construct $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}$. For example, for coalition $\mathcal{S}=\{\text{AI},\text{Systems}\}$, we construct $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}=(36.7,5.4,6.9,10.2)$, enabling a direct feature comparison of the item being explained ($\textbf{v}_{\text{Georgia Tech}}$) against the baseline sample ($\textbf{u}_{\text{Stanford}}$). We wish to understand why Georgia Tech is ranked higher than Stanford (Figure 4c), and why Stanford is ranked higher than UMich (Figure 4d). In both cases, Georgia Tech and UMich have lower values than Stanford for all features except Systems. In Figure 4c, we see that indeed all features except Systems contribute negatively in the relative rank of Georgia Tech as opposed to Stanford. However, the Systems value of Georgia Tech is high enough to rank it higher than Stanford. Reversely, for UMich, we see that, while Systems is the most important feature in the top-10% stratum (Figure 4a), Stanford is ranked higher because of the aggregate contributions of the other features.
The pairwise explanation from ShaRP  is intuitive: for example, Systems improves Georgia Tech’s rank by 5 compared to Stanford. in other words, feature contributions sum to (half) the rank difference between these universities, aligning with Fig. 4a, which highlights Systems as particularly influential for top-$k$ universities.
6. Evaluating ranking explanation methods
We define several metrics for evaluating feature importance methods for ranking. Our focus is on explanation methods that return a numerical vector of feature attributions explaining the outcome for a given item. We can assess pair-wise explanation agreement by comparing the feature vectors of a pair of explanations. Furthermore, we can use an explanation to compute the outcome for the item being explained (e.g., its rank), and compare it to the actual observed outcome for that item. This allows us to assess fidelity of an explanation. Below, we describe explanation agreement and fidelity metrics and also explain how these primitives can be aggregated to assess sensitivity and fidelity of an explanation method, and to quantify inter-method explanation agreement.
Notation
In Section 3, we have been using $\phi(\textbf{v})$ to represent the vector of feature weights, computed using Shapley values. We generalize our definition here to $g(\textbf{v})$ to represent the output of any feature-based explanation method $g$, regardless of whether it consists of Shapley values or of some other numerical quantification of feature importance. For all methods we consider, $g(\textbf{v})$ is a vector of numerical contributions of each feature towards the outcome for item v.
6.1. Fidelity Metrics
Explanation Fidelity
A useful property of feature-based explanations is that the actual outcome can be computed from them. For Shapley-value-based explanations, this follows from the efficiency property of Shapley values, see Section 3. Fidelity measures how well the explanation $g(\textbf{v})$ matches the model prediction $f(\textbf{v})$ being explained, see (Chowdhury et al., 2023; Molnar, 2020). SHAP and LIME explanations can be used to compute an item’s score (score QoI in our terminology), with feature importance indicating the displacement due to that feature from the mean score, either positively or negatively. ShaRP  explanations can be used to compute the outcome for all supported QoIs, including score, rank, and top-$k$, and for the pairwise method.
For the QoIs, namely, score, rank, and top-$k$, we compute fidelity of explanation $g$ for item v as:
(8) $F(g,\textbf{v},\mathsf{QoI}())=1-\frac{1}{Z}\left|{\mathsf{QoI}(\textbf{v})-\sum_{i=1}^{d}g(i,\textbf{v})}\right|$
Here, $\mathsf{QoI}(\textbf{v})$ returns the value of the quantity of interest (i.e., the outcome being explained by $g$), such as v’s score, rank, or presence in the top-$k$, while $g(i,\textbf{v})$ is the contribution of the i-th feature of $v$. Finally, $Z$ is the normalizer set to the maximum distance between a pair of outcomes for the given dataset $\mathcal{D}$ and ranker $f$ (omitted to simplify notation), and for the specified QoI. Note that, for pairwise explanations, fidelity $F(g,\textbf{u}\succ\textbf{v})=1$ if $u$ is ranked higher than $u$ and if $g$ predicts that relative order among the items, and is 00 otherwise.
Example 0.
Consider, for example, the explanation of Texas A&M University’s $rank=34$ in CS Rankings, presented as a waterfall plot in Figure 2b. The sum of feature weights $-20.78-19.52-18.71-1.95=-60.96$ captures the displacement of Texas A&M University in the ranking relative to the middle of the ranked list (position 94.5 out of 189), up to rounding: $94.5-60.96=33.54$. This explanation has near-perfect fidelity $0.9981-\frac{0.16}{189}=$ 0.998. We use the length of the ranked list $Z=189$ as the normalizer for rank QoI.
Method Fidelity
We aggregate per-item fidelity (per Equation 8) to quantify the fidelity of an explanation method as:
(9) $F(g,\mathcal{D})=\mathbb{E}_{\textbf{v}\in\mathcal{D}}F(g,\textbf{v})$
For pairwise, we compute $F(g,\mathcal{D})$ as the expectation of $F(g,\textbf{u}\succ\textbf{v})$ over all pairs of distinct items $\textbf{u},\textbf{v}\in\mathcal{D}$.
6.2. Agreement Metrics
Explanation Agreement
When comparing explanation methods, we may be interested in knowing how similar their explanations are for the same item. Alternatively, when analyzing an explanation method, we may want to know how similar its explanations are for some pair of items (e.g., those that are similar in feature space, or that have similar outcomes, or both).
We define three flavors of explanation agreement, based on (1) Kendall’s tau distance, (2) Jaccard distance of the top-$2$ features, and (3) Euclidean distance between the explanation vectors. For each of these distance metrics, we normalize them to the $[0,1]$ range and then transform their output so that 1 means full agreement (similarity) and 0 means full disagreement. For dataset $\mathcal{D}$ and ranker $f$, we define explanation agreement as:
(10) $A(g,q,\textbf{u},\textbf{v},\mathsf{sim}())=\mathsf{sim}(g(\textbf{u}),q(\textbf{v}))$
[…]
Example 0.
[…]
Method Agreement
To compute agreement for a pair of explanation methods $g$ and $q$, for a dataset $\mathcal{D}$, we compute explanations for each item using each method, compute pair-wise explanation agreement per Eq. 10, and aggregate it across $\mathcal{D}$.
(11) $A(g,q,\mathcal{D},\mathsf{sim}())=\mathbb{E}_{\textbf{v}\in\mathcal{D}}A(g,q,\textbf{v},\textbf{v},\mathsf{sim}())$
Method Sensitivity
The Sensitivity of an explanation method quantifies the similarity between explanations of similar items (Bhatt et al., 2021). We will use $\mathsf{nbr}(\mathbf{v})$ (as in “neighbor”) to refer to a function that retrieves items that are in some sense similar to $\mathbf{v}$, noting that this similarity may be based on items’ features, their outcomes for some QoI, or both. For each v, we retrieve its neighbors $\mathsf{nbr}(\textbf{v})$, compute pair-wise explanation agreement between v and each of its neighbors per Eq. 10, and aggregate this value over $\mathcal{D}$:
(12) $S(g,\mathcal{D},\mathsf{sim},\mathsf{nbr}())=\mathbb{E}_{\textbf{v}\in\mathcal{D},\textbf{u}\in\mathsf{nbr(\textbf{v}})}A(g,g,\textbf{v},\textbf{u},\mathsf{sim}())$
7. The  ShaRP  library
ShaRP  is implemented in Python, follows an API structure similar to scikit-learn (Buitinck et al., 2013), and is parallelized. The library can be used both to compute exact feature importance values and to approximate them to improve running times.
7.1. Implementation of QoIs for ranking
We provide Algorithm 1 to showcase the flexibility of ShaRP. Using this implementation, we can 1) easily switch between QoIs, 2) calculate both the marginal and baseline Shapley values, and 3) approximate the Shapley values for efficiency. The algorithm relies on black-box access to the model that generates the outcome (i.e., specifying an input to the model and observing the outcome used in the QoI). Specifically, algorithm 1 takes as input a dataset $\mathcal{D}$, a reference set $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ from which we draw the samples, an item v for which the explanation is generated, the number of samples to draw $m$, the maximum coalition size $c$, and the iota function $\iota()$ (Equation 5) based on which to quantify feature importance. To change the QoI we change the input $\iota()$ function. To switch to the pairwise baseline method, we set $\mathcal{D}^{\prime}=\textbf{u}$ and $m=1$, where u is the baseline sample we want to compare v to. And finally, to approximate the feature importance, we can use the parameters $m$ and $c$. Passing in the full set of items as the reference set $\mathcal{D}^{\prime}=\mathcal{D}$, and setting $m=|\mathcal{D}-1|$ and $c=|\mathcal{A}-1|$ results in the exact computation of the Shapley values (i.e., the item v’s features are quantified against every other item in $\mathcal{D}$ using all possible coalitions for all features except $i$). Because we are computing the rank of each item against the entire dataset $\mathcal{D}$, we need to provide the dataset in the input along with the reference set. We will provide an empirical analysis of the impact of $m$ and of maximum coalition size on performance in Section 8.3.2.
We now describe the algorithm for the marginal exact computation which encompasses all cases as we explained above. By definition, when computing feature importance, Shapley values use all possible coalitions of features and all items in the dataset. We refer to this as the exact computation of local feature-based explanations. We choose to explicitly provide the construction of the random variable U in lines 4-7 of Algorithm 1 for illustrative purposes. For every feature $i\in\mathcal{A}$, the algorithm computes every coalition of the remaining features $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$. For each coalition $\mathcal{S}$, it draws $m=|\mathcal{D}-1|$ samples from $\mathcal{D}$. Two random vectors of items are created based on these samples: $\textbf{U}_{1}$, in which feature values of the coalition vary as in U and the remaining features are fixed to their values in v; and $\textbf{U}_{2}$, in which features in the coalition and the feature of interest $i$ vary as in U, and the remaining features are fixed to their values of v. The importance of the coalition, $\phi_{i_{\mathcal{S}}}(\textbf{v})$, is computed by invoking the function $\iota()$ that computes the difference in QoI between the two vectors of items $\textbf{U}_{1}$ and $\textbf{U}_{2}$. This value is added to the contribution of feature $i$, $\phi_{i}(\textbf{v})$ weighted by the number of coalitions of this size (${d-1\choose|S|}$, and the total number of possible coalition sizes $d$.
Algorithm 1 Local feature importance using ShaRP
0:  Dataset $\mathcal{D}^{\prime}$, reference set $\mathcal{D}^{\prime}$, item v, number of samples $m$, maximum coalition size c, $\iota()$
1:  $\phi(\textbf{v})=\langle 0,\ldots,0\rangle$
2:  for $i\in\mathcal{A}$ do
3:     for  $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$ and $|\mathcal{S}|\leq c$ do
[…]
5:        $\textbf{U}_{1}=\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$
6:        $\textbf{U}_{2}=\textbf{v}_{\mathcal{A}\setminus\{\mathcal{S}\cup i\}}\textbf{U}_{\mathcal{S}\cup i}$
7:        $\phi_{i_{\mathcal{S}}}(\textbf{v})=\iota(\textbf{U}_{1},\textbf{U}_{2})$
8:        $\phi_{i}(\textbf{v})=\phi_{i}(\textbf{v})+\frac{1}{d}\frac{1}{{d-1\choose|S|}}\phi_{i_{\mathcal{S}}}(\textbf{v})$
9:     end for
10:  end for
11:  return  $\phi(\textbf{v})$, the Shapley values v’s features
Evaluating the $\iota()$ function, is straightforward for the score QoI but not for the ranking-specific QoIs. Specifically, for the score QoI, using the definition in Section 4, we take the mean of the (per-element) difference of $f(\textbf{U}_{1})$ and $f(\textbf{U}_{2})$. However, this is not the case for ranking-specific QoIs. The rank of an item is computed with respect to all other items in the sample. This adds two steps to calculating the rank QoI compared to the score QoI. The item we are explaining needs to be removed from $\mathcal{D^{\prime}}$, and the score of each item $\textbf{u}_{1}\in\textbf{U}_{1}$ (and equivalently $\textbf{u}_{2}\in\textbf{U}_{2}$) needs to be compared to the scores of all items in $\mathcal{D^{\prime}}$. The computation of $\iota_{Rank}$ is summarized in Algorithm 2.
Algorithm 2 $\iota_{Rank}$
0:  Dataset $\mathcal{D}^{\prime}$, scoring function $f$, item v, $\textbf{U}_{1}$, $\textbf{U}_{2}$, number of samples $m$
1:  $\phi=0$
2:  for $i\in\{1,\dots,m\}$ do
3:     $\textbf{u}_{1}=\textbf{U}_{1}(i)$
4:     $\textbf{u}_{2}=\textbf{U}_{2}(i)$
5:     $\mathcal{D}_{1}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{1}\}$
6:     $\mathcal{D}_{2}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{2}\}$
7:     $\phi=\phi+r^{-1}_{\mathcal{D}_{2},f}(\textbf{u}_{2})-r^{-1}_{\mathcal{D}_{1},f}(\textbf{u}_{1})$
8:  end for
9:  return  $\phi/|\textbf{U}_{1}|$
To compute feature importance that explains whether an item appears at the top-$k$, for some given $k$, we use a similar method as for rank QoI. The difference is that rather than computing the difference in rank positions for a given pair of items $\textbf{u}_{1}$ and $\textbf{u}_{2}$, we instead check whether one, both, or neither of them are at the top-$k$. As in Algorithm 2, we work with $\mathcal{D}_{1}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{1}\}$ and $\mathcal{D}_{2}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{2}\}$ for each sample. We increase the contribution to $\phi$ by 1 if only $\textbf{u}_{1}$ is in the top-$k$, and decrease it by 1 if only $\textbf{u}_{2}$ is in the top-$k$. We omit pseudocode due to space constraints.
7.2. Visualizing Feature Importance
We use three methods for visualizing experimental results. The first, used in Figure 2, is a waterfall visualization that presents feature importance for a single item (as in (Lundberg and Lee, 2017)). The second visualization, used in Figures 34a5, and  10, is a box-and-whiskers plot that aggregates local feature importance across ranking strata, each covering 10% of the range. This visualization allows us to see both the median contribution of each feature and the variance in that feature’s contribution across the items in the stratum. The third visualization, used in Figures 4c and 4d, is a bar graph that presents feature contributions for the pairwise method. The visualization is created from the point of view of the first item of the pair.
8. Experimental Evaluation of ShaRP
We conducted numerous experiments to showcase the usefulness and performance of ShaRP  over real and synthetic datasets, with associated score-based ranking tasks. We present efficiency analysis in Section 8.3, conduct a qualitative evaluation of ShaRP  in Section 8.1, and compare it with several other explainability methods in Section 8.2 using the metrics from Section 6. All experiments were conducted on a Linux machine with an Intel Xeon Platinum 8268 CPU 2.90GHz Processor, using 14 cores and 128GB RAM.
We evaluate the performance of ShaRP  and compare it to other local feature importance methods, using several real and synthetic datasets, with the corresponding ranking tasks. Dataset properties, along with ranker type (score-based or learned) are summarized in Table 1 and described below. We present the dataset information and plot the relationship between score and rank for score-based ranking tasks in Appendix A.
Table 1. Datasets used in our experimental study. S stands for score-based ranked task and LtR for learn-to-rank.
name source # tuples # features task
Tennis (ATP) (Joao Fonseca, 2024b) 86 6 S
CS Rankings (CSR) (Berger, 2023) 189 5 S
Times Higher Education (THE) (Joao Fonseca, 2024a) 1,397 5 S
Synthetic (SYN) here 2,000 2 or 3 S
Moving company (MOV) (Yang et al., 2021) 4,000 3 LtR
8.1. Qualitative Analysis
We already presented a detailed case study of CS Rankings presented as an example across the previous sections. To evaluate ShaRP  across different settings, we conducted two additional experiments. First, we analyzed a set of simple synthetic datasets coming from multiple different distributions and studied how each distribution affects the ranking. Secondly, we compared the explanations resulting from two different LtR rankers for the Moving Company dataset.
8.1.1. Score-based Ranking with Synthetic Data
In these experiments presented in more detail in Appendix B, we examine simple datasets consisting of two features and various distributions. Using this simple set-up we examine how different distributions interact with ranking. We use two settings, first, we vary the dataset distributions and fix the scoring function. Secondly, we fix the dataset distributions and vary the scoring functions.
When the scoring function is fixed but the distribution varies, feature importance depends on both the distribution and the stratum. Features with the greatest variance typically have the highest importance for ranking, with items scoring highest on these features ranking at the top. The opposite occurs at the bottom, while in the middle, either feature can dominate, increasing variance in feature importance for both features. For negatively correlated features, the same pattern emerges, but contributions have opposite signs. Discrete features, such as a Bernoulli variable, play a distinct role by segmenting the ranking: items with a value of 1 occupy the top-50%, while those with 0 fall into the bottom-50%. Within each segment, a second feature determines rank.
When we fix the data distribution but vary the ranking function, feature importance can have very different variance per stratum. This depends on the ranking function and on the feature variance. We also show that the feature with the lower variance can become the most important feature when its weight is very large.
Finally, we show that depending on the feature distribution, it is more or less possible for items ranked low to jump to the top-k when they change their values. This contradicts the assumption made in Anahideh and Mohabbati-Kalejahi (2022) where they claim that examining the local neighborhood of an item is sufficient to provide accurate feature importance explanations. We show that even items from the top-50% can move to the top-10% for example under specific distributional assumptions.
(a) XGB (original data)
(b) XGB (fairness interv.)
(c) LGB (original data)
(d) LGB (fairness interv.)
Figure 5. Feature contribution to the rank QoI for (a) XGB over the original moving company dataset, (b) XGB over the unbiased version, (c) LGB over the original moving company dataset, and (d) LGB over the unbiased version.
8.1.2. Learning to Rank
We now showcase how ShaRP  can be used to audit black-box rankers and understand their decision process. We use an XGB ranker with a pairwise ranking objective and an LGB ranker with a LambdaRank objective. Both are trained on training sets and evaluated on test sets of 2,000 tuples each. We use  ShaRP  to explain 100 items (10 per stratum) of each test set, with no approximations and the rank QoI. In Figure 5, we observe that the two LtR models behave significantly differently.
XGB rankers do not appear to rely on the Sex feature, regardless of whether the de-biasing intervention in (Yang et al., 2021) was applied. However, Race plays a significant role; in Figure 5a, Race increases the ranking of all applicants up to the 70th percentile by approximately 400 positions. This is particularly relevant as the Weight Lifting feature contributes positively to ranks in the 70-80th percentile range but negatively in the 60-70th percentile range. Ideally, its importance should be more monotonic, as seen (to some extent) in Figure 5b. While Race has slightly lower relevance overall, it remains an important feature.
In contrast, LGB rankers utilize all features in most cases. Feature importance in an LGB model trained on the original data (Figure 5c) shows that Sex and Race are highly influential across all strata, often ranking as the most important features for applicants in the lowest percentiles (50th onward). Analysis of the 10-20th, 60-70th, and 90-100th percentiles reveals that Weight Lifting has little to no impact on the model’s decisions. Thus, Sex and Race largely determine whether an applicant achieves a high rank. The fairness intervention mitigates this issue to some extent (Figure 5d) by shifting Weight Lifting’s importance toward a more intuitive and desirable role. However, Race and Sex remain highly significant, sometimes even outweighing Weight Lifting.
In summary, results indicate that XGB relies more on Race, while LGB emphasizes Sex. Bias mitigation is effective up to the 10th percentile but fails to correct bias across the remaining strata.
8.2. Comparison to Other Methods
In this section, we compare explainability methods using the metrics from Section 6, focusing on a subset of methods from Section 2. Since ShaRP and Shapley values emphasize individual explanations, we exclude methods designed for global explanations, such as those by Yang et al. and Gale and Marian (Yang et al., 2018; Gale and Marian, 2020). To compare with HIL (Yuan and Dasgupta, 2023), we adapt their code to support any score-based ranker and real data (see Appendix C), and we focus only on their weight-based methods, as their linear-approximation of Shapley values is already covered by SHAP. We do not compare to PrefShap (Hu et al., 2022), which is limited to pairwise datasets using a generalized preferential kernel function.
We compare to HRE (Anahideh and Mohabbati-Kalejahi, 2022) but use only four of their internal methods as provided by their public code base (Decision Trees (DT), Linear Regression (LR), Ordinary Least Squares (OLS), and Partial Least Squares (PLS)) and their default neighborhood settings (5-10 consecutive positions above and below the item being explained). We compare to DEXER which fits a linear regression model to the ranks and explains this model using the score-based SHAP instead of the original blackbox, treating rank as a score. Finally, we compare to SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b), due to their wide use and availability, even though they are not designed for ranking.
8.2.1. Sensitivity
(a) HRE LR
(b) DEXER
(c) LIME
(d) SHAP
(e) ShaRP Rank
(f) HIL Std Rank
Figure 6. Sensitivity results for CS Rankings. Each dot represents a neighbor of the reference item. The x-axis shows the Euclidean explanation distance, the y-axis the absolute rank difference, and the hue indicates feature similarity to the reference item. Methods using rank as the profit function (ShaRP and HIL Std rank) perform better, with ShaRP leading. These are the only methods that consistently provide similar explanations for items that are both feature-similar and have similar outcomes.
Figure 6 presents the results comparing sensitivity for all methods. In this figure, we evaluate the similarity of explanations for pairs of similar items. For each pair of items, we compute three distances: (1) Euclidean distance between the explanations (x-axis); (2) distance between the ranks of the two items (y-axis); and (3) Euclidean distance between the items in terms of their feature values (hue, where lighter means closer). To make the plot, we place one item (the reference item) at position (0,0) and use a scatter point for each other item (neighbor), indicating the distance in ranks and the distance of the explanations. The color of the scatter point indicates the distance between the features of the reference point and the neighbor. We then overlay the plots for all items in the dataset, so that all items are used as reference points and all items are used as neighbors.
Intuitively, items with similar features and close rankings should have similar explanations. That is, the points should align near the diagonal $y=x$, with a darkening hue as the distance from the origin increases. However, this rarely holds in practice. For instance, if one feature dominates, proximity in feature space may not imply explanation similarity. Likewise, items that appear very different in feature space may still receive similar outcomes. Ideally, explanations should reflect these nuances: differing for closely ranked items with distinct features and never being too similar for items far apart in ranking. This suggests that the plot should fill the space below $y=x$, with hue darkening as the distance from the origin increases.
In Figure 6, only methods using the rank QoI produce explanations that follow the expected shape. Both ShaRP  with rank QoI (Figure 6e) and HIL-Std-Rank (Figure 6f, using our expanded definition) successfully generate similar explanations for similarly ranked items with similar features. ShaRP  performs slightly better in clustering points into denser similarity groups. In contrast, SHAP (Figure 6d), a score-based method, relies primarily on feature distances. Its plot shows distinct bands that darken away from 0 and below the diagonal. However, items ranked very differently but with similar features still receive nearly identical explanations.
LIME (Figure 6c), another score-based method, appears to incorporate both feature and ranking distance but fails to distinguish explanations as clearly as rank-based methods.
We see that score-based methods are unable to account for nonlinear relationships between score and rank. DEXER (Figure 6b), which uses linear regression to predict rank and SHAP to explain it, performs almost identically to score-based methods. While a non-linear prediction model might predict ranks more accurately, we follow a more general approach and directly incorporate ranks into the Shapley value utility function.
Finally, the explanation distance of the HRE method (Figure 6a) does not appear to depend on the ranking distance or the feature distance as we see items with similar and dissimilar explanations across all ranks and hues. This is expected, as the explanations depend only on neighbors in terms of ranking (that can have different features) and there is no guarantee that a different cluster of explanations for a further neighborhood will be different from the neighborhood of the reference point.
In Appendix D, we present sensitivity results for all methods, extensively compare ShaRP  with rank QoI to HIL with Std rank, and conduct a similar analysis for a score-based task. We show that ShaRP , which leverages Shapley values, outperforms HIL Std rank, which directly uses model weights, across all datasets. Additionally, we compare ShaRP with score QoI to ShaRP with rank QoI, demonstrating that for score-based tasks, the score QoI aligns with the diagonal line—highlighting the importance of selecting a QoI relevant to the explainability question.
In summary, explanations for the rank QoI, which we are introducing in this paper, are able to more accurately explain ranking tasks compared to other local feature-based explanation methods.
8.2.2. Fidelity
It is possible to calculate Fidelity for SHAP, LIME, ShaRP, and the HIL-score. It is impossible to compute Fidelity for HIL-rank and all the HRE methods. All methods except HIL-score perform very well.
We compute the Fidelity averaged across all items in all datasets. All methods are executed using their recommended settings to compute explanations for score QoI. Additionally, we compute fidelity for ShaRP  for the rank QoI. Recall that ShaRP  is the only method that can compute a Shapley value explanation for this QoI. LIME, SHAP and ShaRP  are all achieving high explanation fidelity, on average ranging from 0.94-0.98, 0.97-1.00, and 1.00 correspondingly. HIL has reasonable fidelity for CSR (0.85) but does not perform consistently on other datasets ranging from 0.14-0.64. We include a table to visualize these results in Appendix D.
8.2.3. Agreement Between Explanations
Figure 7 presents agreement between ShaRP  and all other methods averaged across all datasets. We also provide the agreement per dataset in Appendix D. We use rank and score QoIs for this comparison, as they match those used by the methods we evaluate. Kendall’s tau distance is computed to enable cross-method comparisons. We observe that explanations vary significantly by method, regardless of the QoI. ShaRP aligns most closely with LIME and SHAP across both rank and score QoIs. HRE, which relies on localized information, naturally differs. However, even among HRE variants, explanations remain inconsistent. The two HIL methods and the two ShaRP methods produce similar explanations despite using different QoIs, suggesting that explanation agreement depends on both the QoI and the underlying explanation method In contrast, DEXER, which fits a linear regression to the ranking output and applies SHAP, differs greatly from all methods, indicating that rank cannot be effectively explained without a rank QoI.
Figure 7. Method agreement averaged across all datasets
8.3. Efficiency and Approximation
Table 2. Running time of exact computation.
avg. time (sec)
dataset # tuples # features score rank pair
ATP 86 6 0.004 0.026 0.004
CSR 189 5 0.002 0.022 0.003
THE 1,397 5 0.011 0.423 0.007
SYN 2,000 3 0.002 0.126 0.003
(a) Speedup vs. sample size, THE
(b) Speedup vs. max coalition size, THE
(c) Fidelity vs. max coalition size, THE
Figure 8. Computational time performance of approximation for Times Higher Education (THE), CSRankings (CSR), and ATP Tennis (ATP). Speedup is computed in comparison to exact computation times, reported in Table 2.
8.3.1. Running time of exact computation
In our first experiment, we measure the running time of the exact computation for the rank and score QoIs, and the pairwise method using the rank QoI for 3 of the real and 1 of the synthetic datasets in Table 1. We include only 1 synthetic dataset because all synthetic datasets have the same number of points ($m=2,000$) and at most 3 features, and differ in the structure of the correlations within the data. However, these correlations do not impact the running time of exact computation. Finally, we do not include the running time for the top-$k$ QoI because it is based closely on rank QoI in our implementation, and the running times of generating explanations for these two QoIs are indistinguishable.
Table 2 presents the results, reporting the time to generate an explanation per point, averaged over 100 points for CSR, THE, and SYN, and over 83 points (dataset size) for ATP. We observe that rank QoI is the most challenging to compute, taking 0.423 seconds per item for THE, followed by score (0.011 seconds for THE) and then by pairwise (0.007 seconds for THE).
We observe that the running time of rank and score QoIs increases with both the number of items ($m$ in Algorithm 1) and the number of features $d$. This is expected, as exact feature importance for score and rank QoIs scales linearly with dataset size $\mathcal{D}$ ($m=|\mathcal{D}-1|$ samples for exact computation) and exponentially with $d$ (due to $12^{d}-$ 1 feature coalitions). Pairwise methods, like all baseline methods, use only two data points (the pair), making their running time independent of $m$ but exponential in $d$. However, we use the pairwise method with the rank QoI which requires computing the new rank of an item or pair upon intervention in Shapley value computation (line 7 in Algorithm 2). This rank computation can be implemented in various ways. In our implementation, it scales linearly with $m$, explaining why pairwise QoI for THE ($m=1,397$, $d=5$) has a higher running time than ATP ($m=86$, $d=6$).
8.3.2. Running time and quality of approximation
We implement two methods for approximating feature importance to reduce the running time: reduce the number of samples and restrict maximum coalition size. In addition to measuring the running time, we also measure and report explanation fidelity (Eq. 8) to quantify the quality of the approximation. Figure 8 presents these results for the THE dataset, where exact computation was slowest and so there is most substantial room for improvement. In Appendix D.4 we show similar results for ATP and CSR.
Figure 8a shows the speed-up achieved by reducing the number of samples $m$. Lowering $m$ from 1378 (exact) to 20, while maintaining a maximum of size 4 coalitions (the largest possible for 5 features), accelerates rank QoI by a factor of 32—reducing runtime from 0.423 sec to 0.013 sec. Crucially, this performance gain does not compromise fidelity, which remains above 0.98 (out of 1) across all sample sizes in all experiments.
Figures 8b and 8c show the speed-up and fidelity, respectively, when coalition size is bounded. We observe that speed-up is most substantial for all QoIs for coalitions of size 1, but that fidelity is lower for this approximation method. We observe fidelity values of at least 0.83 for rank, 0.91 for score, and 0.98 for pairwise. Fidelity further improves for rank and score for coalitions of size at most 2, reaching 0.87, 0.94, and 1, respectively. The difference between the approximation methods is expected because they affect different aspects of the sampling process. Sample size affects the number of samples we draw while maximum coalition size affects the number of features we can sample concurrently.
In summary, reducing the number of samples and bounding coalition size are both effective at improving the running time, while retaining high explanation fidelity and method agreement in practice. Nonetheless, unless otherwise stated, we used exact computation in this section to reduce variability in the reported results.
9. Conclusions and Future Work
We introduced ShaRP , a comprehensive framework for quantifying feature importance in selection and ranking. Given the impact of rankers on individuals, organizations, and populations, understanding their decisions is crucial for auditing and compliance (ensuring legal adherence), recourse (helping individuals improve outcomes), and design (optimizing ranking procedures). Our work addresses the interpretability needs of these tasks.
We demonstrated the effectiveness of ShaRP  through a qualitative analysis of an impactful real-world task—the ranking of Computer Science departments. This was complemented by a thorough evaluation on real and synthetic datasets, revealing that our defined quantities of interest provide valuable and complementary insights beyond simple score-rank relationships. We showed that feature importance varies with data distribution even when the scoring function is fixed and exhibits locality. Finally, we compared ShaRP  to other local feature-based explanation methods, showing it performs favorably.
A limitation of our work is that our evaluation of the usefulness of explainability did not include a user study. Understanding how humans make sense of feature importance in ranking and act on their understanding, is in our immediate plans. We also plan to support additional quantities of interest, and develop a comprehensive benchmark that compares both local and global feature importance methods for ranking in terms of their usability, expressiveness, and performance.
Appendix A Datasets
A.1. Datasets
We evaluate the performance of ShaRP  and compare it to other local feature importance methods, using several real and synthetic datasets, with the corresponding ranking tasks. Dataset properties, along with ranker type (score-based or learned) are summarized in Table 1 and described below. We show the relationship between score and rank for score-based ranking tasks in Figure 9.
CS Rankings (CSR) ranks 189 Computer Science departments in the US based on a normalized publication count of the faculty across 4 research areas: AI, Systems (Sys), Theory (Th), and Interdisciplinary (Int) (Venetia Pliatsika, 2023). We use publication data for 2013-2023, with the scoring function provided by CSRankings, a geometric mean of the adjusted counts per area, with # of sub-areas as exponent:
$f=\sqrt[27]{(AC_{AI}^{5}+1)(AC_{Sys}^{12}+1)(AC_{Th}^{3}+1)(AC_{Int}^{7}+1)}$
ATP Tennis (ATP) is based on publicly available 2020-2023 performance data of tennis players from the Association of Tennis Professionals (ATP) (Joao Fonseca, 2024b). We use 2022 data that includes 5 performance-related attributes of 86 players. We select 2022 because this is the year in which data for all 5 attributes is available for the highest number of players. We use the following scoring function that we recovered from the ATP site using the scores:
[…]
Times Higher Education (THE) is a dataset of worldwide university rankings (Joao Fonseca, 2024a). It contains the university name, country, and the scores assigned to the university by Times Higher Education for teaching (TEA), research (RES), citations (CIT), income (INC), and international students (INT). We use 2020 data, for consistency with Anahideh and Mohabbati-Kalejahi (Anahideh and Mohabbati-Kalejahi, 2022) who also used it in their paper, with the scoring function provided by THE:
$f=0.3\times TEA+0.3\times RES+0.3\times CIT+0.025\times INC+0.075\times INT$
(a) Tennis (ATP)
(b) CS Rankings (CSR)
(c) Times Higher Education (THE)
(d) a representative synthetic dataset
(e) Moving Company - train
(f) Moving Company - test
Figure 9. The relationship between an item’s score ($y$-axis) and its rank ($x$-axis) for four score-based tasks.
Moving Company
The moving company scenario (Yang et al., 2021) simulates a hiring process where job applicants are ranked based on their qualification score, computed as a function of their weight lifting ability, sex, and race. We train two different rankers, over two scenarios:
  1. (1)
    Using the original data from a previous hiring process from that company, where female applicants generally display lower weight-lifting ability than male applicants and a lower qualification score. In addition, black applicants have a lower qualification score compared to white applicants, but similar weight-lifting ability. Hence black females face greater discrimination compared to the rest of the applicants.
  2. (2)
    After applying the intersectional fairness intervention proposed in the same paper over the data.
All versions of this dataset (both scenarios and train/test sets) contain 2000 tuples.
We use an eXtreme Gradient Boosting (XGB) and a Light Gradient-Boosting (LGB) Machine to model the rankings of the applicants in the training set, and infer and compute the feature contributions of the applicants in the test set, i.e., all results reported in this section correspond to the test set. The XGB ranker was defined with the pairwise ranking objective, while the LGB ranker uses the lambdarank objective.
Synthetic datasets
We also use numerous synthetic datasets to showcase specific quantitative and qualitative aspects of local feature-based explanations and metrics, and to study specific aspects of performance. These datasets contain 2,000 tuples. In five of them, items have 2 features, $x_{1}$ and $x_{2}$, distributed according to the uniform, Gaussian, or Bernoulli distributions, with varying parameters. We experiment with both independent and correlated features. Each synthetic dataset consists of $2,000$ items. We use three linear scoring functions: $f_{1}=0.8\times x_{1}+0.2\times x_{2}$, $f_{2}=0.5\times x_{1}+0.5\times x_{2}$, and $f_{3}=0.2\times x_{1}+0.8\times x_{2}$.
To explore correlations further, we create three datasets that have three Normal features $x_{1}$, $x_{2}$, and $x_{3}$, and 2,000 items. In the first dataset, all features are independent. In the second, we draw $x_{1}$ and $x_{2}$ from the 2D Gaussian and they are negatively correlated with correlation -0.8. The third feature $x_{3}$ is independent. For the third dataset, we draw the features from the 3D Gaussian. $x_{1}$ and $x_{2}$ are negatively correlated with correlation -0.8, $x_{1}$ and $x_{3}$ are positively correlated with correlation 0.6, and $x_{2}$ and $x_{3}$ are negatively correlated with correlation -0.2. For all three datasets, we use the same scoring function $f_{4}=0.33\times x_{1}+0.33\times x_{2}+0.34\times x_{3}$.
Appendix B Distributional Analysis for Ranking
Fixed scoring function, varying data distribution.
In this experiment, we illustrate that feature importance is impacted by the data distribution of the scoring features to a much greater extent than by the feature weights in the scoring function. Further, we show that feature importance varies by rank stratum. In Figure 10, we show rank QoI for 4 synthetic datasets with the same scoring function $f_{2}$.
We observe that, while the features have equal scoring function weights, their contributions to rank QoI differ for most datasets. In $D_{1}$, the Bernoulli-distributed $x_{2}$ determines whether the item is in the top or the bottom half of the ranking, while the Gaussian-distributed $x_{1}$ is responsible for the ranking inside each half. For $D_{2}$, the uniform $x_{1}$ has higher importance because it often takes on larger values than the Gaussian $x_{2}$. In $D_{4}$, $x_{1}$ and $x_{2}$ are negatively correlated, so when one contributes positively, the other contributes negatively. Only for $D_{3}$, with two uniform identically distributed features, the median contributions of both features are approximately the same within each stratum.
Additionally, we see that feature contributions differ per rank stratum. For example, for $D_{3}$, the medians show a downward trajectory across strata. This is because they quantify the expected change (positive or negative) in the number of rank positions to which the current feature values contribute. Also for $D_{3}$, feature contributions have higher variance in the middle of the range, because a 40-60% rank corresponds to many feature value combinations.
Figure 10. The rank QoI using $f_{2}=0.5\times x_{1}+0.5\times x_{2}$ for four datasets; $D_{1}$: $x_{1}\sim N(0.5,0.1)$, $x_{2}\sim Bern(0.5)$; $D_{2}$: $x_{1}\sim[0,1]$, $x_{2}\sim N(0.5,0.1)$; $D_{3}$: $x_{1}\sim[0,1]$, $x_{2}\sim[0,1]$; $D_{4}$: $x_{1}\sim N(0.5,0.05)$, $x_{2}\sim N(0.75,0.016)$, with -0.8 correlation. Feature contributions are different per rank stratum and data distribution.
Figure 11. Rank QoI for $D_{5}$: $x_{1}\sim N(0.5,0.1)$, $x_{2}\sim N(0.5,0.05)$. Subplots correspond to different scoring functions: $f_{1}=0.8\times x_{1}+0.2\times x_{2}$ (top), $f_{2}=0.5\times x_{1}+0.5\times x_{2}$ (middle), $f_{3}=0.2\times x_{1}+0.8\times x_{2}$ (bottom).
Fixed data distribution, varying scoring function.
In this experiment, we investigate the impact of the scoring function on rank and top-$k$ QoI for two datasets. In Figure 11, we use $D_{3}$ and see that the contributions to rank QoI vary depending on the scoring function. For $f_{1}$, $x_{1}$ is the only important feature (although it carries 0.8 — and not 1.0 — of the weight). This can be explained by the compounding effect of the higher scoring function weight and higher variance of the distribution from which $x_{1}$ is drawn. Between $f_{2}$ and $f_{3}$, features $x_{1}$ and $x_{2}$ switch positions in terms of importance, and show a similar trend, despite being associated with different scoring function weights (0.5 & 0.5 vs. 0.2 & 0.8). This, again, can be explained by the higher variance of $x_{1}$, hence, $x_{2}$ needs a higher scoring function weight to compensate for lower variance and achieve similar importance.
Figure 12. Top-$k$ QoI for $k=10\%$, $D_{2}$: $x_{1}\sim[0,1]$, $x_{2}\sim N(0.5,0.1)$, and $D_{3}$: $x_{1}\sim[0,1]$, $x_{2}\sim[0,1]$. Subplots correspond to different scoring functions: $f_{1}=0.8\times x_{1}+0.2\times x_{2}$ (top), $f_{2}=0.5\times x_{1}+0.5\times x_{2}$ (bottom).
Top-$k$ access.
Access to the top-$k$ is determined by the interaction between the scoring feature weights and the distributions of these features. The top-k QoI tells us how important each feature is when we consider only access to the top-$k$. A positive feature contribution signifies that changing the feature’s value will result in decreased chances of getting to the top-$k$. A very high (or very low) value shows that the changes are significant. Figure 12 illustrates this for datasets $D_{2}$ and $D_{3}$. When we consider two identical uniform features that have equal weights ($D_{2}$ under $f_{2}$), we first notice that their control of top-$k$ access is identical as expected. Additionally, we see that for the top-10, changing either feature would reduce access to the top-k (the values are both very positive). However, for each stratum up to the top-70%, changing either feature can contribute either positively or negatively.
When we consider two identical uniform features ($D_{1}$) one of which has a higher weight ($f_{1}$) or dataset $D_{2}$ (under either $f_{1}$ or $f_{2}$), we see again that for the top-10, changing either feature would reduce access to the top-k. Also, we see that how high the contributions are depends on the distributions. However, we see that for the top 10%-20% changing the second, less important feature would increase the chances of getting into the top-$k$. For the rest of the strata, with some variations depending on the dataset and function, changing the most important feature provides non-zero probability of moving to the top-$k$ and interestingly, this persists even for the lower strata. Evidence that items from lower strata can move to the top-$k$ under some scoring functions and feature distributions counters the assumption of Anahideh and Mohabbati-Kalejahi (2022) that changes in rank are localized.
Appendix C Implementation of HIL
HIL  (Yuan and Dasgupta, 2023) is the only other method that recommends the usage of ranks as a profit function for individual explanations in tabular ranking. While this method is not general, we are interested in comparing it with our rank QoI. This was not straightforward because the method is available as a web app that works only for linear weight scoring functions and datasets of two Gaussian features. To compare the rank-relevance contributions introduced in that paper to the rank QoI, we adapted their method using their definitions and code. This implementation is available alongside our own. Further, we extended their method to work with the specific non-linear scoring function used by CS Ranking, by changing the way that Std rank and Std score (discussed below) are computed.
More specifically, because HIL  (Yuan and Dasgupta, 2023) works only with linear weight scoring functions, they do not provide a full Shapley values implementation but use the linear weights to approximate Shapley values assuming feature independence, see Corollary 1 in (Lundberg and Lee, 2017) and also (Strumbelj and Kononenko, 2014). This is a well-established method to compute Shapley values for linear weights also implemented by SHAP, so we do not compare with this part of the method. In addition, HIL defines two methods to acquire feature contributions: “standardized Shapley values” and “rank relevance Shapley values,” which we will call Std score and Std rank, respectively. Those are not calculated using the linear weight method described above, but rather directly from the weights, and without using the mean score or rank. For an item v, each feature $i$ contribution for Std score is $\phi_{i}=\frac{\beta_{i}\textbf{v}_{i}}{\sum_{\textbf{u}\in\mathcal{D}}f(u)}$, where $\beta_{i}$ is the weight for feature $i$. In other words, the contribution of each feature for each item is the score contribution of this feature over the sum of all scores for all items. Similarly, for Std rank, the contribution of feature $i$ for an item v is $\phi_{i}=\beta_{i}\textbf{v}_{i}\alpha_{\textbf{v}}$, where $\alpha_{\textbf{v}}$ is a scaling factor used to transform the score of the specific item to the rank of the specific item calculated as $\alpha_{\textbf{v}}=\frac{(\max_{r\in r_{\mathcal{D}}}(r)-r^{-1}_{\mathcal{D}(\textbf{v})})\sum_{\textbf{u}\in\mathcal{D}}f(u)}{\max_{r\in r_{\mathcal{D}}}(r)f(\textbf{v})}$. Note that neither of the two formulas is computing Shapley values; rather, they assign a contribution to the features based on the linear weights, and the score and rank. This implies that our rank QoI is the only rank QoI for Shapley values.
Appendix D Additional Details on Method Comparisons
D.1. Fidelity
We provide more details on the Fidelity results discussed in Section 8.2.2. We compute the Fidelity of all the methods that have that property across all datasets. We use SHAP and LIME out-of-the-box so their performance is not perfect (although extremely good). We make this choice to highlight the importance of using exact Shapley values when computing local explanations, where the error in each separate explanation is important as each explanation impacts a separate person.
LIME SHAP ShaRP HIL
dataset score score score rank score
ATP 0.98 1.00 1.00 1.00 0.14
CSR 0.95 0.99 1.00 1.00 0.85
THE 0.94 0.97 1.00 1.00 0.64
Syn 0 0.95 0.99 1.00 1.00 0.37
Syn 1 0.95 0.99 1.00 1.00 0.29
Syn 2 0.95 0.99 1.00 1.00 0.35
Table 3. Fidelity across all methods across all datasets.
D.2. Agreement between Explanations
Figure 13 provides a per-dataset visualization of the agreement between the explanations of the methods in Section 8.2.3.
In Fig. 13a We visualize Kendall’s tau explanation distance correlation of ShaRP using the rank QoI with all other methods across every dataset. In Fig. 13b, we plot the same result for ShaRP using the score QoI. As in the aggregated method agreement plot (Fig. 7), ShaRP is very similar to SHAP and LIME for both QoIs. As expected it is more similar to the score QoI but not identical that is perhaps because we used SHAP out-of-the-box which applies some approximation parameters for running time optimization. Similarly, ShaRP behaves similarly to what we discussed in Fig. 7 to all methods across the datasets except the HIL methods for the Synthetic datasets. We hypothesize that this is because the HIL methods are able to perform better for those datasets due to their Synthetic nature.
(a) Method Agreement between Sharp using the rank QoI and all methods across all datasets
(b) Method Agreement between Sharp using the score QoI and all methods across all datasets
Figure 13. Method Agreement
D.3. Sensitivity
We provide the sensitivity analysis for all methods for CSRankings in Fig. 14. In addition, to HRE LR, DEXER, LIME, SHAP, ShaRP rank, and HIL std rank that we presented in Fig. 6, we also plot HRE DT in Fig. 14a, HRE OLS in Fig. 14c, HRE PLS in Fig. 14d, ShaRP score in Fig. 14g, HIL std score in Fig. 14i. We see that all HRE methods perform similarly or worse than HRE-LR, this is unsurprising as all these methods are used locally. We also see that both HIL std score and ShaRP score perform similarly to SHAP, this is also expected. HIL std score and DEXER are very similar which reveals our inability to predict the rank using the ranked output of the model. Specifically, HIL std score assumes knowledge of the weights used by the model and uses them directly to compute the feature importance. DEXER is assuming black-box access to the ranked output only and fits a linear regression model on the ranking. Nevertheless, judging from these results, it appears that DEXER is explaining the score (and not the rank) and is learning the model weights to do so. Inadvertently, we also show that the choice of the explanation method makes a big difference to the final explanation.
(a) HRE DT
(b) HRE LR
(c) HRE OLS
(d) HRE PLS
(e) LIME
(f) SHAP
(g) ShaRP Score
(h) ShaRP Rank
(i) HIL Std Score
(j) HIL Std Rank
(k) DEXER
Figure 14. Sensitivity analysis for the CSRankings dataset for all methods
To further compare ShaRP  with rank QoI and HIL with Std rank, we present Figure 15. Even though both methods are appropriate for the ranking task we are examining, in this figure, we see that ShaRP  with rank QoI (Figures 15a,  15c, and  15e) can capture the full range of different ranks and features, and that groups the items more successfully. HIL with Std rank cannot capture the difference of feature values for ATP (Figure 15b), or the similarly ranked items that have different feature values for the Synthetic experiment (the middle area close to the $x$-axis of Figure 15f). Both methods perform similarly for THE (Figures 15c and  15d).
(a) ShaRP Rank ATP
(b) HIL Std Rank ATP
(c) ShaRP Rank THE
(d) HIL Std Rank THE
(e) ShaRP Rank Synthetic-0
(f) HIL Std Rank Synthetic-0
Figure 15. Sensitivity analysis for the ATP, THE, and Synthetic dataset 0 for the methods using the rank QoI.
Finally, we present an analysis of ShaRP using the score QoI and the rank QoI for the CSR dataset but for a score task (instead of rank). The goal of this analysis is to show that the sensitivity of the methods that use a score QoI is very high when we are explaining a score task. In other words, if we are trying to explain the score, then the methods that use a score-based profit function perform the best as it is fully expected.
The task we are going to explain is the score of the CSRankings scoring function. We choose this task for two reasons, first, we already provided the results of the CSRankings ranking task and we can draw a direct comparison. Secondly, we have a ranking for that dataset and we can plot the methods that use the rank QoI for juxtaposition. Note that it is entirely redundant to use a rank-based QoI method in this case. In fact, it is redundant to even produce a ranking as we are asking an explainability question about the score. But we are choosing to provide this information to showcase that each explainability task needs its own profit function and the choice of the profit function makes a big difference to the final explanation.
In Fig. 16, we evaluate the similarity of explanations for pairs of similar items when we attempt to explain the score. For each pair of items, we compute three distances: (1) Euclidean distance between the explanations (x-axis); (2) distance between the scores (instead of rank) of the two items (y-axis); and (3) Euclidean distance between the items in terms of their feature values (hue, where lighter means closer). To make the plot, we place one item (the reference item) at position (0,0) and use a scatter point for each other item (neighbor), indicating the distance in ranks and the distance of the explanations. The color of the scatter point indicates the distance between the features of the reference point and the neighbor. We then overlay the plots for all items in the dataset, so that all items are used as reference points.
Unlike Fig. 14, we now expect to see items that are both similar in terms of their features and scored near each other to have similar explanations. We would still expect all points to be on or near the diagonal line $y=x$, with the hue getting darker as we move away from the origin, if their explanations successfully explain the score.
In Fig. 16, we see that indeed the score-based methods have the desired shape we discussed in Section 8.2.1. ShaRP score is extremely similar and almost entirely fits the $y=x$ line. ShaRP rank, appears to be providing explanations that do not depend on the score distance between the items’ outcomes (y-axis) or the feature distance between the items (hue) as expected.
This analysis shows how QoI selection is important when providing an explanation. The score is unable to perform well for a ranking task since it estimates the impact of each feature on the score outcome and similarly, it is completely unreasonable to use a rank QoI when explaining the score.
(a) ShaRP Score
(b) ShaRP Rank
Figure 16. Sensitivity analysis for the CSRankings dataset for all methods when the task we are trying to explain is a score task. Compared to Figure 14 we see that the methods that are using a score QoI are performing better.
(a) Speedup vs. sample size, CSR
(b) Speedup vs. sample size, ATP
(c) Speedup vs. maximum coalition size, CSR
(d) Speedup vs. maximum coalition size, ATP
Figure 17. Computational time performance of approximation for CSRankings (CSR), and ATP Tennis (ATP). Speedup is computed in comparison to exact computation times, reported in Table 2.
(a) Fidelity vs. maximum sample size, THE
(b) Fidelity vs. maximum sample size, CSR
(c) Fidelity vs. maximum sample size, ATP
(d) Fidelity vs. maximum coalition size, CSR
(e) Fidelity vs. maximum coalition size, ATP
Figure 18. Fidelity of approximation for Times Higher Education (THE), CSRankings (CSR), and ATP Tennis (ATP), using different sample sizes and maximum coalition sizes.
D.4. Additional Efficiency and Approximation Additional Results
In this section, we present the extended results previewed in Section 8.3.2.
In Figure 17 we present the speedup vs. sample size, and speed-up vs max coalition size for THE, CSR, and ATP. We already presented the results for THE in Figure 8. We observe similar results but scaled down due to the dataset sizes. In Figure 18, we present the corresponding fidelity for both sample size and max coalition size. In this plot we also provide the fidelity of THE for sample size that we omitted in the main body of this work due to space constraints. We observe that fidelity is very high for all sample sizes, and almost identical to the fidelity of THE for all max coalition sizes.
In Figures 19 and 20, we present the method agreement between the approximation and the exact computation for CSRankings (CSR). We omit method agreement results for the other datasets, where ShaRP  performs similarly. In 19a and 19b, we present agreement of the approximation when we vary sample size for the rank and the score QoI. We evaluate the agreement using the Jaccard Index (considering the top-2 features), Kendall’s tau distance, and the Euclidean distance of the feature vectors (converted to unit vectors). Here, we see that performance is similar for both QoIs. The Jaccard index is over 0.9 for any sample size, and is the distance metric with the worst performance for both QoIs. This is worth noting as shorter explanations are often considered more interpretable (Molnar, 2020). Agreement is similar or higher for all QoIs when we vary maximum coalition size, see Figure 20a20c.
(a) Agreement of ShaRP across different sample sizes using maximum coalition size for rank QoI
(b) Agreement of ShaRP across different sample sizes using maximum coalition size for score QoI
Figure 19. Agreement of ShaRP for CSRankings when varying the sample size.
(a) Agreement of ShaRP across different coalitions sizes using maximum sample size for rank QoI
(b) Agreement of ShaRP across different coalitions sizes using maximum sample size for score QoI
(c) Agreement of ShaRP across different coalitions sizes using maximum sample size for pairwise QoI
Figure 20. Agreement of ShaRP for CSRankings when varying the coalition size.